### v1

# Separating subsets from their images

Abstract.
Let $G$ be a transitive permutation group. In this paper, we introduce and study the parameter $\mathbf{m}(G)$, which estimates the size of the largest set of points $A$ such that there exists a permutation $g$ with the property that $A\cap A^{g}$ is empty. In particular, we focus on deriving general bounds for arbitrary transitive groups, and on the asymptotic behaviour of certain families of primitive groups.
Key words and phrases:
self-separable; regular; primitive
2020 Mathematics Subject Classification:
Primary: 20B05; Secondary: 05B30, 20B15, 20B25
The research presented in this paper was initiated during a research visit by MB and KR to IMFM, funded by Slovenian Research and Innovation Agency, research programme number P1-0294. They gratefully acknowledge the support and hospitality provided. MB is supported by the Slovenian Research Agency program P1-0222 and grant J1-50001, and he is a member of the Italian GNSAGA INdAM research group. ML and $P$ are supported by the Slovenian Research and Innovation Agency, programme number P1-0294.
1. Introduction
Let us begin by presenting a graph-theoretical problem [^1] We thank Urban Jezernik for pointing this problem to us. which spurred our interest in the topic of the paper and led us to define and investigate the notions presented here.
Suppose we are given a simple finite graph $\Gamma$ on $n$ vertices and we want to know whether its complement contains an isomorphic copy of $\Gamma$ as a subgraph, or equivalently, whether two copies of $\Gamma$ can be packed into the complete graph $\mathbf{K}_{n}$. Clearly, if $\Gamma$ contains more than $n\choose 2$} edges, then such a packing is impossible. On the other hand, if $\Gamma$ is very sparse (for example, if it has maximum valence $1$ and at least $3$ vertices), then the packing will very likely exist. It is thus natural to ask what is the largest integer $m$ such that every graph on $n$ vertices and at most $m$ edges can be packed twice into $\mathbf{K}_{n}$. Since such a packing is clearly impossible for the complete bipartite graph $\mathbf{K}_{1,n-1}$, we see that $m\leq n-2$. Surprisingly, as was ingeniously proved in [12, Theorem 1], this bound is sharp, in the sense that every graph on $n$ vertices and at most n2n-2 edges lies in its complement as a subgraph.
What we just discussed for graphs can actually be translated into the language of permutation groups. Suppose $\Gamma$ is a graph with vertex-set $n]:=\{1,2,\ldots,n\}$ and let $E\subseteq{[n]\choose 2}$ be its edge-set. If there exists an isomorphism from $\Gamma$ to a graph contained in the complement of $\Gamma$, then this isomorphism is a permutation of $[n]$ which, when viewed as a permutation on the unordered pairs $n]\choose 2$}, maps $E$ to a set disjoint from $E$. In short, two copies of $\Gamma$ can be packed into $\mathbf{K}_{n}$ if, and only if, the group $\mathrm{Sym}(n)$ in its action on $n]\choose 2$} contains an element $g$ such that $E\cap E^{g}=\emptyset$; we shall say that such a set $E$ is self-separable—see Definition 1. The graph-theoretical theorem mentioned in the previous paragraph can now be translated into the assertion that the size of a smallest non-self-separable set for the action of $\mathrm{Sym}(n)$ on $n]\choose 2$} is n1n-1.
As we will try to show in this paper, putting the original packing problem in this general group-theoretical context reveals connections to several classical problems and theorems in combinatorics, finite geometry and group theory, as well as opens a plethora of new and interesting questions, some of which will be addressed in this paper.
Definition 1.
Let $G$ be a transitive permutation group on a finite set $\Omega$ with $|\Omega|\geq 2$. A subset $A\subseteq\Omega$ is said to be self-separable for $G$ provided that there exists $g\in G$ such that $A\cap A^{g}=\emptyset$, and is non-self-separable otherwise. Further, let
$\mathbf{m}(G):=\min\left\{|A|\in\mathbb{N}\mid A\subseteq\Omega\hbox{ is not self-separable for }G\right\}$
be the size of a smallest non-self-separable set for $G$.
Observe that the parameter $\mathbf{m}(G)$ is also equal to the largest integer such that all sets of size less than that integer are self-separable, that is:
$\mathbf{m}(G)=\max\left\{t\in\mathbb{N}\mid\forall A\subseteq\Omega:|A|<t\Rightarrow A\hbox{ is self-separable}\right\}.$
The main theme of this paper is to address the following:
Problem A.
Given a transitive permutation group $G$, determine the parameter $\mathbf{m}(G)$.
The above question can of course be approached from several angles, ranging from algorithmic considerations to exact theoretical results.
The computational task of determining $\mathbf{m}(G)$ for a given permutation group $G$ appears to be highly challenging. The only algorithm we devised during the preparation of this paper involves an exhaustive search (see [5] and Section 7): Systematically testing all subsets $A$ of the domain whose sizes exceed the theoretical lower bound we will establish in C, and stopping when we find the minimal example $A$ that is not self-separable for $G$. When $G$ is known to be $k$-homogenous, we may limit the search space to subset containing a fixed $k$-subset. No further improvements of this naïve approach are known to us. Therefore we ask the following:
Question 2.
Does there exist a significantly more efficient algorithm for computing the parameter $\mathbf{m}(G)$ given a permutation group $G$ (perhaps, subexponential in the degree)?
In this paper, however, we shall be especially interested in asymptotic (lower and upper) bounds on $\mathbf{m}(G)$ in terms of the degree of $G$. As we shall see, for every group $G$ of order $n$, the parameter $\mathbf{m}(G)$ always satisfies $\sqrt{n}\leq\mathbf{m}(G)\leq\lceil(n+1)/2\rceil$ (see C and Example 5). Moreover, infinite families of transitive groups exist for which $\mathbf{m}(G)$ is asymptotically square root of the degree (for example, the family of symmetric groups acting on unordered pairs) as well as infinite families where $\mathbf{m}(G)$ is asymptotically linear in the degree of the group (for example, the symmetric groups in their natural action). What we find interesting, though, is that there seems to be a dichotomy between these square root and linear asymptotical behaviours.
Let us present the remaing players of this paper. For an infinite family $\mathcal{C}$ of finite permutation groups and for an integer $n$, let $\mathcal{C}_{n}$ be the subfamily of permutation groups in $\mathcal{C}$ that have degree $n$, and let
$X_{\mathcal{C}}:=\left\{n\in\mathbb{N}\mid\mathcal{C}_{n}\neq\emptyset\right\}.$
The asymptotic behaviour of the parameter $\mathbf{m}(G)$ for $G\in\mathcal{C}$ will be analysed in terms of the functions:
$\mathbf{f}_{\mathcal{C}}:X_{\mathcal{C}}\to\mathbb{N},\quad n\mapsto\min\left\{\mathbf{m}(G)\mid G\in\mathcal{C}_{n}\right\},$
and
$\mathbf{F}_{\mathcal{C}}:X_{\mathcal{C}}\to\mathbb{N},\quad n\mapsto\max\left\{\mathbf{m}(G)\mid G\in\mathcal{C}_{n}\right\}.$
Remark 3.
We can immediately observe that, for every two families $\mathcal{D}\subseteq\mathcal{C}$ and for every positive integer $n\in X_{\mathcal{C}}$,
$\mathbf{f}_{\mathcal{C}}(n)\leq\mathbf{f}_{\mathcal{D}}(n)\leq\mathbf{F}_{\mathcal{D}}(n)\leq\mathbf{F}_{\mathcal{C}}(n).$
To understand the behavior of these functions on large class of transitive permutation groups, we shall tackle the following problem.
Question 4.
Given an infinite class $\mathcal{C}$ of transitive permutation groups, what is the asymptotic behaviour of $\mathbf{f}_{\mathcal{C}}(n)$ and of $\mathbf{F}_{\mathcal{C}}(n)$? In particular, do there exist two positive constants $r\in\left(\frac{1}{2},1\right)$ and $c$, and an infinite family $\mathcal{C}$ such that
$\liminf_{n}\frac{\mathbf{f}_{\mathcal{C}}(n)}{n^{r}}=\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}}(n)}{n^{r}}=c?$
That is, does there exist a family $\mathcal{C}$ such that the expected asymptotic behaviour of $\mathbf{m}(G)$, for a generic $G\in\mathcal{C}$, is neither square root nor linear?
To give you the taste of the statements we will be proving, let us list some of them. In the case of the class $\mathcal{C}_{{\rm tr}}$ of all transitive permutation groups, we prove that $\liminf\mathbf{f}_{\mathcal{C}_{{\rm tr}}}(n)$ grows asymptotically as $\sqrt{n}$ (see Theorem D), while $\limsup\mathbf{F}_{\mathcal{C}_{{\rm tr}}}(n)$ is asymptotically $\frac{n}{2}$ (see Theorem C), or to be precise:
$\liminf_{n}\frac{\mathbf{f}_{\mathcal{C}_{{\rm tr}}}(n)}{\sqrt{n}}=1\quad\hbox{ and }\quad\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}_{{\rm tr}}}(n)}{n}=\frac{1}{2}.$
Further, as is stated in Theorem E, for the class $\mathcal{C}_{{\rm reg}}$ of regular permutation groups, the asymptotic bound for $\mathbf{F}_{\mathcal{C}_{{\rm reg}}}$ can be significantly improved to
$\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}_{{\rm reg}}}(n)}{\sqrt{n}}\leq\frac{4}{\sqrt{3}}.$
A particular attention is also given to the class $\mathcal{C}_{{\rm pr}}$ of primitive permutation groups, where we start to address the following problem.
Problem B.
Determine what is the largest subclass $\mathcal{C}\subseteq\mathcal{C}_{{\rm pr}}$ where a bound of the form
$\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}}(n)}{\sqrt{n}}\leq C$
holds for some constant $C$.
To facilitate smooth reading, the next section provides an extended abstract of the results proved in this paper. The proofs can be found in later sections.
2. Results
2.1. General bounds
We start our investigation by determining some general bounds on $\mathbf{m}(G)$.
Theorem C.
Let $G$ be a transitive permutation group of degree $n$, and let $\alpha$ be a point. Then
(A) $\mathbf{m}(G)\geq\frac{1}{2|G_{\alpha}|}+\sqrt{n-\frac{1}{|G_{\alpha}|}+\frac{1}{4|G_{\alpha}|^{2}}},$
with equality if, and only if, $G$ is a regular group of collineations of a projective plane. In particular, if $\mathcal{C_{\mathrm{tr}}}$ is the class of all transitive permutation groups, then
$\liminf_{n}\frac{\mathbf{f}_{\mathcal{C_{\mathrm{tr}}}}(n)}{\sqrt{n}}=1.$
The proof of C (given in section Section 4) relies on the Neumann Separation Lemma (see [8, 37]), and the characterisation of the equality depends on considerations from incidence geometry.
We now focus on an upper bound for $\mathbf{m}(G)$. Clearly, for every transitive permutation group $G$ of degree $n$, two subsets containing strictly more than $\lfloor n/2\rfloor$ points cannot have trivial intersection. Hence,
(B) $\mathbf{m}(G)\leq\left\lceil\frac{n+1}{2}\right\rceil.$
The following example shows that the bound is sharp.
Example 5.
Let $G$ be either the symmetric group $\mathrm{Sym}(n)$ or the alternating group $\mathrm{Alt}(n)$ endowed with their natural actions on $[n]$. Since $G$ is $\lfloor n/2\rfloor$-transitive, every subset containing half of the points or less can be mapped to its complement by a permutation in $G$. Hence, every subset $A$ containing $\lfloor n/2\rfloor$ points is self-separable for $G$. Therefore,
$\mathbf{m}(\mathrm{Sym}(n))=\mathbf{m}(\mathrm{Alt}(n))=\left\lfloor\frac{n}{2}\right\rfloor+1=\left\lceil\frac{n+1}{2}\right\rceil.$
Therefore, we obtain the following asymptotic behavior.
Theorem D.
Let $\mathcal{C}_{\mathrm{tr}}$ be the class of all transitive permutation groups. Then
$\lim_{n}\frac{\mathbf{F}_{\mathcal{C_{\mathrm{tr}}}}(n)}{n}=\frac{1}{2}.$
Classifying the groups that attain the upper bound in Inequality (B) is a nontrivial problem, posing challenges both computationally and theoretically. We provide this classification in Section 7. Broadly speaking, unless the degree of $G$ is less than $24$, $G$ is either $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$ in their natural actions on $n$ points, or it is isomorphic to a subgroup of $C_{2}\mathop{\mathrm{wr}}\mathrm{Sym}(m)$ or $\mathrm{Sym}(m)\mathop{\mathrm{wr}}C_{2}$ endowed with the imprimitive action on $2m$ points (where the permutation group that $G$ induces on the system of blocks or on a single block, respectively, is either alternating or symmetric). For a precise statement, including the complete list of small-degree groups meeting the bound, we refer to Lemmas 21 and 22.
2.2. Permutation groups with bounded stabiliser
In the previous section, we determined the asymptotic behaviour of $\mathbf{f}_{\mathcal{C}_{\mathrm{tr}}}$ and $\mathbf{F}_{\mathcal{C}_{\mathrm{tr}}}$ for the class of all transitive permutation groups. Note however, that the gap between the square root asymptotics of $\liminf\mathbf{f}_{\mathcal{C}_{\mathrm{tr}}}(n)$ and the linear growth of $\mathbf{F}_{\mathcal{C}_{\mathrm{tr}}}(n)$ leaves many questions about the parameter $\mathbf{m}(G)$ for a generic transitive permutation group $G$ open. In what follows, we will try to identify subclasses $\mathcal{C}\subseteq\mathcal{C}_{\mathrm{tr}}$ for which we are able to exert a more effective control. That is, families for which the gap between $f_{\mathcal{C}_{\mathrm{tr}}}$ and $F_{\mathcal{C}_{\mathrm{tr}}}$ is smaller. In fact, we will observe that for many important families of permutation group $\mathcal{C}\subseteq\mathcal{C}_{\mathrm{tr}}$, $\mathbf{F}_{\mathcal{C}}(n)$ is asymptotically proportional to $\sqrt{n}$ (as well as $\mathbf{f}_{\mathcal{C}}(n)$, by consequence). The class of regular permutation groups provides a first example supporting this viewpoint.
Theorem E.
Let $G$ be a regular permutation group of degree $n$. Then
$\mathbf{m}(G)\leq\frac{4}{\sqrt{3}}\sqrt{n}.$
In particular, if $\mathcal{C}_{\mathrm{reg}}$ is the family of all regular permutation groups, then
$1=\liminf_{n}\frac{\mathbf{f}_{\mathcal{C}_{\mathrm{reg}}}(n)}{\sqrt{n}}\leq\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}_{{\mathrm{reg}}}}(n)}{\sqrt{n}}\leq\frac{4}{\sqrt{3}}.$
This result is essentially proven in Remark 18, where it is shown that, for a subset $A$ not being self-separable for a regular group $G$ is equivalent to $A$ being a difference basis for $G$:
Definition 6.
A subset $A$ of a group $G$ such that $AA^{-1}=G$ is a difference basis for $G$
Let us remark that the asymptotics for the minimal size of a difference basis for a generic group $G$ is still an open problem in additive combinatorics.
If every nontrivial element of $G$ can be written uniquely as $ab^{-1}$ for some $a,b\in A$, then the difference basis $A$ is called a difference set, which are a widely used tool in finite geometries. If $q$ is the order of a classical projective plane, the celebrated paper [45] by Singer proves the existence of a difference set of size $q+1$ for the cyclic group of order $q^{2}+q+1$. Recall that these permutation groups are precisely those that meet the lower bound in C.
[…]
Corollary F.
For every positive integer $d$, there exists a constant $\mathbf{c}(d)$ such that, for every primitive permutation group $G$ of degree $n$ and minimal nontrivial subdegree at most $d$, the following inequality holds
$\mathbf{m}(G)\leq\mathbf{c}(d)\sqrt{n}.$
In particular, if $\mathcal{C}$ is the family of primitive permutation groups of minimal nontrivial subdegree at most $d$, then
$\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}}(n)}{\sqrt{n}}\leq\mathbf{c}(d).$
Corollary G.
For every positive integer $d$, there exists a constant $\mathbf{c}(d)$ such that, for every $2$-arc-transitive group of automorphisms $G$ of a connected graph with $n$ vertices and of valency at most $d$, the following inequality holds
$\mathbf{m}(G)\leq\mathbf{c}(d)\sqrt{n}.$
In particular, if $\mathcal{C}$ is the family of $2$-arc-transitive group of automorphisms of connected graphs of valency at most $d$, then
$\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}}(n)}{\sqrt{n}}\leq\mathbf{c}(d).$
2.3. Reduction to primitive permutation groups
A standard approach in permutation group theory is to investigate whether and to what extent the problem can be reduced to the setting of primitive groups.
Recall that every imprimitive permutation group $G$ can be reduced to an iterated wreath product of primitive permutation groups as follows. Let $\Sigma$ be a system of imprimitivity for $G$ and let $B$ be a block in $\Sigma$. As is standard, we denote by $G_{B}^{B}$ the permutation group that the setwise stabiliser $G_{B}$ induces on $B$, and by $G^{\Sigma}$ the permutation group that $G$ induces on $\Sigma$. Then $G$ embeds into the imprimitive wreath product $G_{B}^{B}\mathop{\mathrm{wr}}G^{\Sigma}$. Moreover, if $\Sigma$ is a coarsest system of imprimitivity, then $G^{\Sigma}$ is primitive, meanwhile if $\Sigma$ is a finest system of imprimitivity, then $G_{B}^{B}$ is primitive.
(The following theorem is proved in Section 3.2.)
Theorem H.
Let $G$ be a group, and suppose that $G$ embeds into the imprimitive wreath product $K\mathop{\mathrm{wr}}H$. Then
$\mathbf{m}(K)+\mathbf{m}(H)-1\leq\mathbf{m}(G)\leq\mathbf{m}(K)\mathbf{m}(H)\,.$
The upper bound is asymptotically sharp. Let $K$ and $H$ be isomorphic to the cyclic group of degree $q^{2}+q+1$ with regular action, and let $G=K\mathop{\mathrm{wr}}H$. Note that the degree of $G$ is $(q^{2}+q+1)^{2}$. For large $q$, using C both to find the lower bound for $\mathbf{m}(G)$ and the upper bounds for $\mathbf{m}(K)$ and $\mathbf{m}(H)$,
$q^{2}+o(q^{2})\leq\mathbf{m}(G)\leq\mathbf{m}(K)\mathbf{m}(H)\leq(q+o(q))^{2}=q^{2}+o(q^{2}).$
This proves the asymptotic sharpness of the upper bound of H. The same reasoning, applied with $K$ and $H$ taken as regular groups and using E, shows that $\mathbf{m}(G)\leq\frac{8}{3}\sqrt{n}$ – matching the optimal lower bound up to a multiplicative constant.
On the other hand, we are unable to prove the optimality of the lower bound. Even showing the existence of a family of permutation groups for which the lower bound is asymptotically tight up to a constant requires further work. The following ingredient, of independent interest, will be essential. (Its proof is given in Section 6.)
Theorem I.
If $G$ is the wreath product $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$ endowed with the imprimitive action on $ab$ points, then
[…]
In particular, if $G$ is an imprimitive permutation group with a block system of cardinality $b$ whose blocks contain $a$ points, then
$\mathbf{m}(G)\leq a+b-1\,.$
Let $a$ and $b$ be two even positive integers, and let $K=\mathrm{Sym}(a)$ and $H=\mathrm{Sym}(b)$. By choosing $G=K\mathop{\mathrm{wr}}H$, I implies
$\mathbf{m}(G)=a+b-1=2\mathbf{m}(K)+2\mathbf{m}(H)-5.$
On the other hand, H yields
$\mathbf{m}(G)\geq\mathbf{m}(K)+\mathbf{m}(H)-1.$
If we allow $a$ and $b$ to grow as two arbitrary unbounded functions, we find that, asymptotically, $\mathbf{m}(G)$ is twice the lower bound of H – proving the asymptotic sharpness up to the multiplicative constant $2$. This leads to the following question.
Question 7.
Is the lower bound stated in H asymptotically sharp? If it is not, is $2(\mathbf{m}(K)+\mathbf{m}(H))$ an asymptotically sharp lower bound? Furthermore, by excluding an appropriate infinite family of transitive groups, is it possible to obtain a larger lower bound?
2.4. Primitive permutation groups
Recall that, by Example 5, for $\mathrm{Alt}(n)$ and $\mathrm{Sym}(n)$ in their action on $n$ points, $\mathbf{m}(\mathrm{Alt}(n))=\mathbf{m}(\mathrm{Sym}(n))=\lceil(n+1)/2\rceil$. At the moment, they constitute the the single infinite family of primitive permutation groups where we can prove that a linear asymptotic behavior is exhibited. We suspect that, possibly apart from few well-understood infinite families, the size of the smallest non-self-separable subset is asymptotically proportional with the square root of the degree of the group. This intuition is captured by the following question.
Question 8.
Does there exist a constant $C$ such that, for the infinite class $\mathcal{C}_{{\rm pr}}^{\ast}$ of all primitive permutation groups apart from alternating and symmetric groups in their natural action,
$\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}_{{\rm pr}}^{\ast}}(n)}{\sqrt{n}}=C?$
Natural candidates that might exhibit a different behaviour are primitive permutation groups of product action type whose socle is a direct product of alternating groups in their natural action (see M and Lemma 21). If the answer to 8 has a negative answer, then we still propose B, already stated in Section 1, as a possible direction of investigation.
In view of this and the already established square-root lower bound (see C), we focus on obtaining upper bounds for $\mathbf{m}(G)$, and thus dismissing the computation of $\mathbf{f}_{{\mathcal{C}_{{\rm pr}}^{\ast}}}(n)$. Let us start our investigation with an example.
Example 9.
Let $\mathrm{PSL}(d,q)\unlhd G\leq\mathrm{P}\Gamma\mathrm{L}(d,q)$ act on the projective space $\mathrm{PG}(d-1,q)$. To find an upper bound on $\mathbf{m}(G)$, one needs to find a subset $A$ of the projective space such that every image of $A$ under a collineation always intersects $A$. An obvious candidate is a projective subspace of dimension exceeding half of that of the ambient $\mathrm{PG}(d-1,q)$. In this case, we can check that if d1d-1 is even, we can take a subspace of dimension (d1)/2(d-1)/2, and we find that, for large $n$,
$\mathbf{m}(G)\leq\sqrt{n}+o(\sqrt{n}).$
Otherwise, if d1d-1 is odd, we take a subspace of dimension $d/2$, and hence we find
$\mathbf{m}(G)\leq n^{\frac{d}{2(d-1)}}+o\left(n^{\frac{d}{2(d-1)}}\right).$
This construction is not meaningful for $d=2$, for which we cannot take any projective subspace apart from the line itself. In fact, we know very little of the behaviour of $\mathbf{m}(G)$, where $G\leq\mathrm{Aut}(PG(1,q))$. Furthermore, even for $d\geq 4$, the exponent d/2(d1)d/2(d-1) is always larger than $1/2$ (though it converges to $1/2$ as $d$ goes to infinity), perhaps suggesting that these groups might also provide a negative answer to 8. However, we believe that subsets distinct from subspaces might provide an example of smaller non-self separable set. It would be fascinating if some subgeometry $A$ of $\mathrm{PG}(d-1,q)$ were not self-separable, but all the obvious candidates (such as ovoids or Galois subgeometries) seem to be self-separable or too large.
Similar reasoning can produce a large array of upper bounds for primitive groups of almost simple type in standard action. As in [30], with the term standard actions, we refer to permutation groups that are either alternating, symmetric or classical, and, in the former two cases, the domain consists of $k$-subsets, while in the latter the groups are endowed with a subspace action. (For the list of subspace actions, we refer to [11, Chapter 4].)
[…]
$(a)$ $\mathrm{Alt}(m)$ $k$-subsets $\lesssim\frac{k!^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}}{\left\lceil\frac{k}{2}\right\rceil!}n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(b)$ $\mathrm{PSL}_{d}(q)$ $k$-subspaces $\lesssim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(c)$ $\mathrm{PSL}_{d}(q)$ pairs $(X,Y)$ of subspaces in direct sum, with $\dim(X)=k$ $\lesssim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(d)$ $\mathrm{PSL}_{d}(q)$ pairs $(X,Y)$ of subspaces with $X\leq Y$, $\dim(X)=k$, $\dim(Y)=d-k$ $\lesssim n^{\frac{k\left\lceil\frac{d-3k}{2}\right\rceil+(d-k)\left\lceil\frac{k}{2}\right\rceil}{k(2d-3k)}}$
$(e)$ $\mathrm{PSU}_{d}(q)$ totally isotropic $k$-subspaces $(\clubsuit)$
$(f)$ $\mathrm{PSU}_{d}(q)$ nondegenerate $k$-subspaces $\lesssim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(g)$ $\mathrm{PSp}_{2d}(q)$ totally isotropic $k$-subspaces $(\clubsuit)$
$(h)$ $\mathrm{PSp}_{2d}(q)$ nondegenerate $2k$-subspaces $\lesssim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(i)$ $\mathrm{P}\Omega_{2d}^{+}(q)$ totally singular $k$-subspaces $(\clubsuit)$
$(j)$ $\mathrm{P}\Omega_{2d}^{-}(q)$ totally singular $k$-subspaces $(\clubsuit)$
$(k)$ $\Omega_{2d+1}(q)$ totally singular $k$-subspaces, $q$ odd $(\clubsuit)$
$(l)$ $\Omega_{2d}^{\epsilon}(q)$ nonsingular $1$-subspaces, $q$ even $(\heartsuit)$
$(m)$ $\mathrm{P}\Omega_{2d}^{+}(q)$ nondegenerate hyperbolic $2k$-subspaces $\lesssim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(n)$ $\mathrm{P}\Omega_{2d}^{+}(q)$ nondegenerate parabolic $(2k+1)$-subspaces, $q$ odd $(\spadesuit)$
$(o)$ $\mathrm{P}\Omega_{2d}^{+}(q)$ nondegenerate elliptic $2k$-subspaces $\lesssim n^{\frac{1}{k}\left\lceil\frac{k+1}{2}\right\rceil}$
$(p)$ $\mathrm{P}\Omega_{2d}^{-}(q)$ nondegenerate parabolic $(2k+1)$-subspaces, $q$ odd $(\spadesuit)$
$(q)$ $\mathrm{P}\Omega_{2d}^{-}(q)$ nondegenerate elliptic $2k$-subspaces $\lesssim n^{\frac{1}{k}\left\lceil\frac{k+1}{2}\right\rceil}$
$(r)$ $\Omega_{2d+1}(q)$ nondegenerate hyperbolic $2k$-subspaces, $q$ odd $\lesssim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(s)$ $\Omega_{2d+1}(q)$ nondegenerate elliptic $2k$-subspaces, $q$ odd $\lesssim n^{\frac{1}{k}\left\lceil\frac{k+1}{2}\right\rceil}$
Table 1. Upper bound for the standard actions of the primitive groups of almost simple type. The asymptotic notation and the symbols $(\spadesuit)$, $(\clubsuit)$ and $(\heartsuit)$ are explained in Remark 10.
Theorem J.
Let $G$ be a primitive group of almost simple type in standard action on the domain $\Omega$. Then, an asymptotic upper bound for $\mathbf{m}(G)$ can be found in the third column of Table 1.
Remark 10.
If $f,g:\mathbb{N}\to\mathbb{R}^{+}$ are two functions such that the limit of the quotient $\limsup f(n)/g(n)$ lies in the interval $[0,1]$, then we write $f\lesssim g$. Hence the expression in the last column of Table 1 consist of the leading term of the asymptotic expansion of the cardinality of the non-self-separable subsets we construct in the proofs.
Observe that in the last column of Table 1 the symbols $(\spadesuit)$, $(\clubsuit)$ and $(\heartsuit)$ appear. The first symbol $(\spadesuit)$ appears every time the socle of the classical group into consideration is orthogonal and the action is on nondegenerate parabolic $(2k+1)$-subspaces. In this case, we have a precise value for the asymptotic upper bound computed, which can be read from Table 4. The symbol $(\clubsuit)$ appears whenever we encounter the action of a classical group on totally isotropic or totally singular subspaces. Although we have an explicit upper bound for $\mathbf{m}(G)$, which can be derived from Table 3, the expression are so complicated that we cannot access a precise asymptotic, even if aided by a calculator. Still, some experimentation with these functions shows that, if one fixes $k$ and allows $d$ to grow to infinity, the expected outcome is the asymptotic behavior $\lesssim\sqrt{n}$. Last, the symbol $(\heartsuit)$ appears in case $(l)$: if $k=1$, in all the other cases our method gives trivially an upper bound, and we have not investigated this case by itself to reach a different conclusion.
Remark 11.
For most standard actions, for a fixed even value of the parameter $k$ in Table 1, we see that $\mathbf{m}(G)\lesssim\sqrt{n}$. Hence, if the degree is large enough, we can expect $\mathbf{m}(G)$ to be arbitrarily close to the lower bound given in C. Moreover, if $k$ is odd, we can note that, as $k$ gets larger, the upper bound computed in J gets arbitrarily close to $\sqrt{n}$.
Let now move on to the other types of primitive groups as described in the O’Nan–Scott Theorem [29]. In what follows, we adopt the partition into eight classes proposed in [38]. We will recall the properties of the classes we address in Section 8.
We start with primitive permutation groups of simple diagonal type.
Theorem K.
Let $G$ be a quasiprimitive permutation group of simple diagonal type, let $k\geq 3$ be the number of direct simple factors of its socle, and let $\varepsilon$ be $1$ if the socle type is a group of Lie type, and 0 otherwise. Then,
$\mathbf{m}(G)\leq 4\left(\frac{16}{3}\right)^{\frac{k-1}{2}}n^{\frac{1}{2}+\frac{1}{k-1}}\left(\frac{1}{4k}\log_{2}n\right)^{\epsilon}.$
In particular, if $\mathcal{C}$ is the family of all quasiprimitive groups of simple diagonal type, then
$\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}}(n)}{n^{\frac{1}{2}+\frac{1}{k-1}}\log_{2}(n)}\leq\frac{1}{k}\left(\frac{16}{3}\right)^{\frac{k-1}{2}}.$
The proof of K, provided in Section 8.2, also holds for primitive permutation groups of type holomorph of simple and holomorph of compound, but in these cases the resulting upper bound for $\mathbf{m}(G)$ is meaningless as it exceeds $n$. We leave all the holomorph types out of our analysis.
We now concentrate on primitive groups of type product action and of type compound diagonal. They can be described as being the output of the construction of a primitive wreath product applied to types almost simple and simple diagonal. This fact inspired us to obtain the following general theorem (proved in Section 3.2).
Theorem L.
Let $G=H\mathop{\mathrm{wr}}K$ be a permutation group in product action on the domain $\Omega=\mathrm{Fun}(\Gamma,\Delta)$. Then,
$\mathbf{m}(G)\leq\mathbf{m}(H)^{|\Gamma|}\,.$
Therefore, we obtain immediately these corollaries.
Corollary M.
Suppose that there exists a constant $C$ such that 8 has a positive solution for the class $\mathcal{C}_{\rm as}^{\ast}$ of all the primitive permutation groups of type almost simple of non alternating socle. Then, for the infinite class $\mathcal{C}_{\rm pa}^{\ast}$ of all primitive permutation groups of type product action whose socle is not isomorphic to a direct product of alternating groups,
$\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}_{\rm pa}^{\ast}}(n)}{\sqrt{n}}\leq C.$
Corollary N.
Suppose that there exists a constant $C$ such that 8 has a positive solution for the class $\mathcal{C}_{\rm sd}$ of all the primitive permutation groups of type simple diagonal with constant $C$. Then, for the infinite class $\mathcal{C}_{\rm cd}$ of all primitive permutation groups of type compound diagonal,
$\limsup_{n}\frac{\mathbf{F}_{\mathcal{C}_{\rm cd}}(n)}{\sqrt{n}}\leq C.$
This concludes the extended summary of our work, while the rest of the paper is devoted to the proofs.
3. Stability of $\mathbf{m}(G)$
In this section, we investigate the stability of $\mathbf{m}(G)$, $\mathbf{f}_{\mathcal{C}}(n)$ and $\mathbf{F}_{\mathcal{C}}(n)$ under certain natural operations applied either to the group $G$ or to the family $\mathcal{C}$.
3.1. Stability with respect to subgroups and subfamilies
Recall that $\mathcal{C}$ is an infinite family of transitive permutation group, and that $\mathcal{C}_{n}$ is the subfamily of permutation groups of degree $n$.
Observe that, for every $G\in\mathcal{C}_{n}$, every subset $A$ of the domain of $G$ satisfying $|A|<\mathbf{f}_{\mathcal{C}}(n)\leq\mathbf{m}(G)$ is self-separable for $G$ by definition. Therefore, to find a lower bound $a(n)+1$ for $\mathbf{f}_{\mathcal{C}}(n)$, we need to prove that, for all $G\in\mathcal{C}_{n}$ and for every subset $A$ in the domain of $G$ such that $|A|\leq a(n)$, $A$ is self-separable for $G$. On the other hand, every subset $A$ of the domain of $G$ with the property that $|A|\geq\mathbf{F}_{\mathcal{C}}(n)\geq\mathbf{m}(G)$ is not self-separable. Hence, to find an upper bound $b(n)+1$ for $\mathbf{F}_{\mathcal{C}}(n)$, we need to prove that, for all $G\in\mathcal{C}_{n}$, there exists a subset $A$ in the domain of $G$ such that $|A|\geq b(n)$ and $A$ is not self-separable for $G$, that is, for every $g\in G$, we can prove that $|A\cap A^{g}|\geq 1$.
Lemma 12.
Let $H\leq G$ be two transitive permutation groups. Then
$\mathbf{m}(H)\leq\mathbf{m}(G).$
Proof.
Let $A$ be a set such that $|A|\leq\mathbf{m}(H)$. By definition of $\mathbf{m}(H)$, there exists a permutation $h\in H$ such that $A\cap A^{h}$ is empty. Since $h\in G$, the bound follows. ∎
3.2. Stability with respect to quotients and extensions
[…]
$\Omega/N=\left\{\omega^{N}\mid\omega\in\Omega\right\}.$
Lemma 13.
Let $G$ be a transitive permutation group on $\Omega$, and let $N$ be a normal subgroup of $G$. Then
[…]
Proof.
[…]
Since there is no clear notion of an extension for permutation groups, to mirror Lemma 13, we focus on examining how the parameter of interest relates between a wreath product $H\mathop{\mathrm{wr}}K$ and its constituents $H$ and $K$. We start with the imprimitive action for $H\mathop{\mathrm{wr}}K$. Explicitly, if $\Delta$ is the domain of $H$ and $\Gamma$ of $K$, $H\mathop{\mathrm{wr}}K$ acts on the Cartesian product $\Delta\times\Gamma$ by the law
$(\delta,\gamma)^{(h_{1},h_{2},\dots,h_{|\Gamma|})k}=(\delta^{h_{\gamma}},\gamma^{k}),$
where $\delta\in\Delta$, $\gamma\in\Gamma$, $(h_{1},h_{2},\dots,h_{|\Gamma|})\in\mathrm{Fun}(\Gamma,H)$ and $k\in K$. (Recall that $\mathrm{Fun}(\Gamma,H)$ is the group of functions from $\Gamma$ to $H$, and it can be identified with the $|\Gamma|$-fold direct product $H^{|\Gamma|}$.)
Proof of H.
Let $G$ be a transitive permutation group that embeds in $H\mathop{\mathrm{wr}}K$ with imprimitive action.
We start by proving the lower bound. Let $A$ be any set of points of size $\mathbf{m}(H)+\mathbf{m}(K)-2$. Either, there exists a block $B$ which contains at least $\mathbf{m}(H)$ points or not. If the latter holds, then, there exists an element $g$ of the base group $\mathrm{Fun}(\Gamma,H)$, such that $A$ and $A^{g}$ intersect trivially. Otherwise, $A$ has a nontrivial intersection with at most $\mathbf{m}(K)-1$ blocks, hence there exists an element $g\in K$ from the top groups that maps the blocks intersecting $A$ to blocks that do not. In any case, we are able to find a permutation $g$ such that $A\cap A^{g}$ is empty, and hence every set of size $\mathbf{m}(H)+\mathbf{m}(K)-2$ is self-separable for $G$. The lower bound immediately follows.
For the upper bound, consider a subset of points
$A=\{(x,y)\mid 1\leq x\leq\mathbf{m}(H),\,1\leq y\leq\mathbf{m}(K)\}\,$,
where we identified the domain of $G$ with the Cartesian product $\Delta\times\Gamma$. By the definition of $\mathbf{m}(H)$ and $\mathbf{m}(K)$, it follows that, for every $g\in G$, $|A\cap A^{g}|\geq 1$. This is enough to prove the upper bound. ∎
Now, we focus on wreath products in product action. The wreath product of $H\mathop{\mathrm{wr}}K$ can be endowed with an action on $\mathrm{Fun}(\Gamma,\Delta)$, which we understand as the $|\Gamma|$-fold Cartesian power of $\Delta$. Indeed, for every $g=(h_{1},h_{2},\ldots,h_{|\Gamma|})k\in H\mathop{\mathrm{wr}}K$, for every $f\in\mathrm{Fun}(\Gamma,\Delta)$ and for every $\gamma\in\Gamma$, we let
$f^{g}(\gamma)=f\left(k^{-1}\gamma\right)^{h_{k^{-1}\gamma}}.$
(For ease of reading, we are denoting the action of $K$ on $\Gamma$ as a left multiplication.)
Proof of L.
Let $R\subseteq\Delta$ be a subset such that, for every $h\in H$, $R\cap R^{h}$ is not empty, and consider
$A=\mathrm{Fun}(\Gamma,R)=\left\{f\in\mathrm{Fun}(\Gamma,\Delta)\mid\forall\gamma\in\Gamma:f(\gamma)\in R\right\}.$
For every $g=(h_{1},h_{2},\ldots,h_{|\Gamma|})k\in H\mathop{\mathrm{wr}}K$, for every $f\in A$ and for every $\gamma\in\Gamma$, we compute
$f^{g}(\gamma)=f\left(k^{-1}\gamma\right)^{h_{k^{-1}\gamma}}\in R^{h_{\gamma k^{-1}}}.$
We now observe that
$A\cap A^{g}=\mathrm{Fun}(\Gamma,R)\cap\left(R^{h_{1k^{-1}}}\times R^{h_{2k^{-1}}}\times\ldots\times R^{h_{|\Gamma|k^{-1}}}\right)$
is not empty due to the choice of $R$. In particular, by applying Lemma 12 and by choosing $R$ of cardinality $\mathbf{m}(H)$, we obtain that
$\mathbf{m}(G)\leq\mathbf{m}(H)^{|\Gamma|}.\qed$
3.3. Stability with respect to change of action
We complete the section by establishing a connection between the sought-after parameter for two different actions of the same abstract groups.
Lemma 14.
Let $G$ be an abstract group, and let $H$ and $K$ be two core-free subgroups of $G$ with $H\leq K$. Then
[…]
Proof.
[…]
$A=\{Ha_{1},Ha_{2},\dots,Ha_{m}\}.$
Without loss of generality, we can rearrange the elements of $A$ so that
$A^{\prime}=\{Ka_{1},Ka_{2},\dots,Ka_{k}\},$
which is a $k$-subset of $G/K$, for some $k\leq m$. By our choice of $m$, there exists a group element $g\in G$ such that $A^{\prime}\cap A^{\prime}g$ is empty, or, more explicitely, such that
$\left(\bigcup_{i=1}^{k}Ka_{i}g\right)\cap\left(\bigcup_{j=1}^{k}Ka_{j}\right)\quad\hbox{is empty}\,.$
Since, for every $i^{\prime}\in\{1,2,\dots,m\}$, there exists $i\in\{1,2,\dots,k\}$ such that $Ha_{i^{\prime}}\subset Ka_{i}$, and since the right multiplication by $g$ is a bijection on $G/H$, it follows that
$\left(\bigcup_{i=1}^{m}Ha_{i}g\right)\cap\left(\bigcup_{j=1}^{m}Ha_{j}\right)\quad\hbox{is empty}\,.$
In particular,
[…]
$B=\{Kb_{1},Kb_{2},\dots,Kb_{k}\}.$
Let $\{k_{1},k_{2},\dots,k_{|K:H|}\}$ be a set of representative for the right cosets of $H$ in $K$, and consider
$B^{\prime}=\{Kb_{1},Kb_{2},\dots,Kk_{1}b_{m},Hk_{2}b_{1},\dots,Hk_{|K:H|}b_{m}\}.$
Observe that $|B^{\prime}|\geq k|K:H|$, and that
$\bigcup_{i=1}^{k}Kb_{i}=\bigcup_{i=1}^{k}\left(\bigcup_{j=1}^{|K:H|}Hk_{j}b_{i}\right).$
Since $B$ is not self-separable for $G$, the previous equality implies that $B^{\prime}$ is not self-separable for $G$. Therefore,
[…]
By taking the minimum on both sides, we obtain the desired lower bound.
Hence, the proof of Lemma 14 is complete. ∎
We would like to point out that Lemma 14 is quite weak in a general setting. For instance, consider the symmetric group $\mathrm{Sym}(n)$ in its natural action (that is, with stabiliser $\mathrm{Sym}(n-1)$) and in its regular action (that is, with a trivial stabiliser). Then, applying the upper bounds from Lemmas 14 and E, we deduce that
[…]
On the other hand, if the index $|K:H|$ can be bounded by a constant, Lemma 14 can produce significant estimates. (For two functions $f,g:\mathbb{N}\to\mathbb{R}^{+}$, we write $f\asymp g$ if there exists a positive integer $M$ and a positive constant $C$ such that, for every $n\geq M$, their quotient $f(n)/g(n)$ is positive and bounded from above by $C$.)
Lemma 15.
Let $\mathcal{C}$ and $\mathcal{D}$ be two families of permutation groups such that, for every abstract group $G$, if $G$ appears in $\mathcal{C}$ with stabiliser $K$, then $G$ appears in $\mathcal{D}$ with stabiliser $H$. Suppose that there is a constant $d$ such that, for every $G$, $|K:H|\leq d$. Then
$\mathbf{f}_{\mathcal{C}}\asymp\mathbf{f}_{\mathcal{D}}\quad\hbox{and}\quad\mathbf{F}_{\mathcal{C}}\asymp\mathbf{F}_{\mathcal{D}}.$
Proof.
The claimed results are obtained simply by taking the minimum and the maximum of the inequalities from Lemma 14 and observing that a linear dilatation of the argument of the function does not change the asymptotic behaviour. ∎
4. Proof of C
We give a proof of Neumann’s separation lemma (see [8, 37] for the seminal version of this result, and [1, Theorem 5.3] for our version), and we derive the lower bound in C as a corollary.
Lemma 16.
Let $G$ be a transitive permutation group on $\Omega$, and let $A$ and $B$ be two subsets of $\Omega$. Then
$\frac{1}{|G|}\sum_{g\in G}|A\cap B^{g}|=\frac{|A||B|}{|\Omega|}.$
Proof.
We define the set
$X=\left\{(\alpha,\beta,g)\in\Omega\times\Omega\times G\mid\alpha^{g}=\beta\right\}\,.$
For every $\alpha,\beta\in\Omega$, the set of elements $g\in G$ satisfying $\alpha g=\beta$ is a right coset of the stabiliser $G_{\alpha}$. Therefore,
(C) $|X|=|A||B||G_{\alpha}|=\frac{|A||B||G|}{|\Omega|}\,.$
On the other hand, for every $g$, we can choose $(\alpha,\beta)$ to complete the triplet in $|A\cap B^{g}|$ ways. Hence,
(D) $|X|=\sum_{g\in G}|A\cap B^{g}|\,.$
The result is obtained by equating Equations C and D. ∎
Proof of C.
Let $G$ be a transitive permutation group on $\Omega$ of degree $n$, and let $A$ be a subset of $\Omega$. By Lemma 16, we have
(E) $\frac{1}{|G|}\sum_{g\in G}|A\cap A^{g}|=\frac{|A|^{2}}{n}\,.$
If there does not exist $g\in G$ such that $A\cap A^{g}$ is empty, then
$\frac{1}{|G|}\sum_{g\in G}|A\cap A^{g}|\geq\frac{1}{|G|}\sum_{g\in G\setminus\{1\}}1+\frac{|A|}{|G|}=1+\frac{|A|-1}{|G|},$
where equality is attained if, and only if, $|A\cap A^{g}|=1$ for every nontrivial $g\in G$.
By Equation E, we also see that
$\frac{|A|^{2}}{n}\geq 1+\frac{|A|-1}{|G|},$
and thus
$|A|\geq\frac{1+\sqrt{1-4|G|(1-|G|)/n}}{2|G|/n}=\frac{1}{2|G_{\alpha}|}+\sqrt{n-\frac{1}{|G_{\alpha}|}+\frac{1}{4|G_{\alpha}|^{2}}}\,.$
Let us denote the expression on the right-hand side by $\alpha(G)$. We have shown that, if a subset $A$ is not self-separable for $G$, its size needs to be at least $\alpha(G)$. Therefore, $\mathbf{m}(G)\geq\alpha(G)$, as required.
Now, we shall prove that the lower bound is met if, and only if, $\Omega$ is the point-set of a projective plane and $G$ is a regular group on $\Omega$. Let
$A^{G}=\{A^{g}\mid g\in G\},$
and consider the incidence structure (or design) $(A^{G},\Omega)$, where $A^{g}$ is incident with $\omega$ if $\omega\in A^{g}$. Observe that $G$ preserves this incidence relation.
Recall that, if $|A|=\alpha(G)$, then $|A\cap A^{g}|=1$ for every nontrivial permutation $g\in G$. Thus, the set-wise stabiliser $G_{A}$ of $A$ is trivial, and hence the action of $G$ is regular on $A^{G}$. The transitivity of $G$ on $\Omega$ and $A^{G}$ gives us that $(A^{G},\Omega)$ is uniform and regular. Clearly, the design is incomplete, and the hypothesis that, for $g\in G\setminus\{1\}$ $|A\cap A^{g}|=1$ implies that it is balanced. Therefore, $(A^{G},\Omega)$ is a balanced incomplete block design, and we can apply Fisher’s inequality. Doing so, we obtain
$|G|=|A^{G}|\leq|\Omega|=n.$
Therefore, $G$ is regular on $\Omega$, and hence the design is symmetric. Using for the last time our hypothesis on the intersections in $A^{G}$, $(A^{G},\Omega)$ is a projective plane, and so is its dual $(\Omega,A^{G})$. ∎
We would like to note here that, if $G$ acts regularly on a non-degenerate projective plane of order $q$, then the right-hand side of Inequality (A) in C is an integer. In fact,
$|A|=\frac{1}{2}+\sqrt{q^{2}+q+1-1+\frac{1}{4}}=\frac{1}{2}+\sqrt{\left(q+\frac{1}{2}\right)^{2}}=q+1.$
Furthermore, we observe that the equality can hold if, and only if, such right-hand side is an integer, which in turn happens if, and only if, $n$ is of the form $n=(m|\alpha|)^{2}+(2|G_{\alpha}|-1)m+1$ for some positive integer $m$. When the right-hand side is not an integer, then the lower bound for $\mathbf{m}(G)$ can be improved by taking the ceiling of that value. This makes us pose the following:
Question 17.
Does there exist a transitive permutation group $G$ of degree $n$ not of the form $(m|G_{\alpha}|)^{2}+(2|G_{\alpha}|-1)m+1$ for any $m\in\mathbb{N}$ such that
$\mathbf{m}(G)=\left\lceil\left(\,\frac{1}{2|G_{\alpha}|}+\sqrt{n-\frac{1}{|G_{\alpha}|}+\frac{1}{4|G_{\alpha}|^{2}}}\,\right)\right\rceil?$
If not, how can the bound in C be improved in this case?
5. Proof of E
We start this section by focusing our attention on groups acting regularly. Then, we will generalize our result to families of permutation group with stabilisers of bounded order.
Remark 18.
Let $G$ be a regular permutation group of degree $n$ on $\Omega$. As usual, we identify $G$ with $\Omega$, so that the action we are dealing with is the action of $G$ on itself by right multiplication. Let $A$ be a subset of $G$. Suppose that there is a $g\in G$ such that $A\cap Ag$ is empty. Then $g$ is not an element of $AA^{-1}$, and thus $AA^{-1}$ is a proper subset of $G$. On the other hand, if $A\cap Ag$ is nonempty for every $g\in G$, then $AA^{-1}=G$. Hence, $\mathbf{m}(G)$ is the size of the smallest set $A$ such that $AA^{-1}=G$.
As per 6, a subset $A$ of a group $G$ such that $AA^{-1}=G$ is called a difference basis for $G$. A closely related notion appears in additive combinatorics, where Rédei and Rényi introduced in [39] the concept of covering the finite subset $\{0,1,2,\dots,n\}$ of the integers by a difference set AAA-A, with $A\subseteq\mathbb{Z}$. The problem of constructing such sets $A$ of minimal possible size remains a relevant question in additive number theory (see, for instance, [6, 43]).
These ideas also intersect naturally with classical topics in combinatorics and finite geometry. For instance, planar difference sets are difference bases $A\subseteq G$ with the additional property that, for every $g\in G$, there exists a unique pair $(a,b)\in A^{2}$ such that $g=ab^{-1}$. Such sets play a fundamental role in the construction of finite projective planes. In particular, in his seminal work [45], Singer showed that every finite projective plane admits a cyclic collineation group of order equal to the number of its points. Oberve that this fact is directly connected to the sharpness of the lower bound in C.
Another related concept is that of a basis for a group $G$: a subset $A\subseteq G$ such that $A^{2}=G$. If $A$ is inverse-closed, then the notions of basis and difference basis coincide. This concept originates with Rohrbach [40], and has since been developed further in [7, 26, 36]. We would like to note that, as observed in [31], the existence of small bases is related to the degree-diameter problem for Cayley graphs of diameter $2$.
The cardinality of a minimal difference basis is bounded below by the size of a minimal difference set, and an upper bound of $3\sqrt{|G|}$ is given in [27]. Determining the exact minimal size of a difference basis remains an active area of research, even for specific families of groups (see [2, 3, 4]). Any improvement to these bounds would lead to immediate improvements in the asymptotic estimates computed in E. For example, for any finite Abelian $p$-group $G$ in its regular right action, where $p$ is a prime with $p\geq 11$, the result of [4] implies that
$\mathbf{m}(G)<\frac{\sqrt{2}(\sqrt{p}-1)}{\sqrt{p}-3}\sqrt{|G|}.$
Proof of E.
Let $G$ be a finite group, and suppose that $G$ acts regularly on itself by right multiplication. By Remark 18, we need to build a difference basis $A$ for $G$ of as small size as possible. By [26, Theorem 1], there exist two positive constants $c_{1}$ and $c_{2}$ and two subsets $B$ and $C$ such that
$|B|=c_{1}\sqrt{|G|},\quad|C|=c_{2}\sqrt{|G|},\quad c_{1}+c_{2}\leq 4/\sqrt{3},\quad BC=G.$
Hence, the subset $A=B\cup C^{-1}$ is a difference basis containing at most $4\sqrt{|G|}/\sqrt{3}$ elements. Indeed,
$AA^{-1}\subseteq G=BC\subseteq(B\cup C^{-1})(B^{-1}\cup C)=AA^{-1}.$
Therefore,
$\mathbf{m}(G)\leq\frac{4}{\sqrt{3}}\sqrt{|G|}.$
The computation of the limit follows using C for $\mathbf{f}_{C}(n)$ and the inequality just proved for $\mathbf{F}_{\mathcal{C}}(n)$. ∎
Surprisingly, [26, Theorem 1] relies on the fact that every group which is not cyclic of prime order contains a subgroup $H$ whose order exceeds $|G|^{1/2}$ (see [28] for a proof of this fact), which in turn relies on the classification of finite simple groups. In a classification free context, if $G$ have even order, the best result known is that there exists a subgroup $H$ with $|H|\geq|G|^{1/3}$, as the famous result of Brauer and Fawler in [10] states. On the other hand, if $G$ is of odd order, then, by the Feit–Thompson Theorem [19], $G$ is solvable, and solvability is enough to prove the existence of a subgroup of order exceeding the square root of the order of the group.
As the constant $4/\sqrt{3}$ in [26, Theorem 1] does not depend on the classification of finite simple groups, E remains valid for solvable groups. This prompts us to ask the following question.
Question 19.
Without using the classification of finite simple groups, can we find a function $\mathbf{g}:\mathbb{N}\to\mathbb{R}$ such that, for every nonsolvable regular permutation groups $G$ of order $n$,
$\mathbf{m}(G)\leq\mathbf{g}(n)?$
We now turn our attention to family of transitive permutation groups whose stabilisers are bounded.
Remark 20.
Observe that, by applying Lemma 15 with $\mathcal{D}$ equal to the family of regular permutation groups, we obtain that, for every $\mathcal{C}$ family of transitive permutation groups with stabilisers of bounded order,
$\mathbf{f}_{\mathcal{C}}\asymp\mathbf{F}_{\mathcal{C}}\asymp\sqrt{n}.$
Actually, since we have an explicit upper bound for $\mathbf{f}_{D}$ and $\mathbf{F}_{\mathcal{D}}$, an immediate computation shows that, if $\mathbf{c}$ is the maximum of the order of a stabiliser in $\mathcal{C}$, then
$\sqrt{n}\leq\mathbf{f}_{\mathcal{C}}(n)\leq\mathbf{F}_{\mathcal{C}}(n)\leq 4\sqrt{\frac{\mathbf{c}n}{3}}.$
We do not put any focus on this precise upper bound because, for the following application, the precise value of $\mathbf{c}$ is not known.
Proof of F.
We want to apply Lemma 15 to the families
[…]
By the positive solution to Sims’s Conjecture in [14], there exists a function $\mathbf{g}$ such that the order of any stabiliser of a permutation group in $\mathcal{C}$ is at most $\mathbf{g}(d)$. Therefore, combining Theorems C and E with Remark 20, we find that there exists a constant $\mathbf{c}(d)>0$ such that, for every
$\sqrt{n}\leq\mathbf{f}_{\mathcal{C}}(n)\leq\mathbf{F}_{\mathcal{C}}(n)\leq\mathbf{c}(d)\sqrt{n}.$
The claims of F immediately follow. ∎
The proof of G follows verbatim that of F, where $\mathcal{C}$ is the family of $G$ which are $2$-arc-transitive groups of automorphisms of some graphs $\Gamma$, whose valency do not exceed $d$, and the positive solution to Sims’s Conjecture is substituted with the positive solution to Weiss’s Conjecture for $2$-arc-transitive graphs obtained in [20, 47, 48, 50, 51].
This result can be actually generalized to a larger class of automorphism groups of graphs. Given a graph $\Gamma$ and a vertex-transitive group of automorphisms $G$, the local group $L$ of the pair $(\Gamma,G)$ is the permutation group that a vertex-stabiliser induces on the neighbourhood of its fixed point. A permutation group $L$ is graph-restrictive if there exists a constant such that, for every pair $(\Gamma,G)$ that witness $L$ as a local group, then the order of a vertex-stabiliser is bounded by this constant. For instance, the property we have used above is that $2$-transitive permutation groups are graph-restrictive. Therefore, our proof actually extends beyond the scope of G. For instance, in [16], it is conjectured that any semiprimitive permutation group is graph-restrictive. For the curious reader, examples of graph-restrictive permutation groups can be found in [21, 22, 33, 34, 46, 49].
6. Proof of I
As stated in the introduction, to discuss the sharpness of H we need to understand $\mathbf{m}(\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b))$, where the wreath product of symmetric groups is endowed with its imprimitive action.
Proof of I.
Let $G=\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$, and let $\Sigma$ be its block system. We denote by $B_{1},\dots,B_{b}$ the blocks of $\Sigma$. Preliminary, we show that $\mathbf{m}(G)\leq a+b-1$. Consider any subset of the domain of the form
$A=B_{1}\cup\{\omega_{2}\}\cup\dots\cup\{\omega_{a}\},$
where $\omega_{i}\in B_{i}$ for $2\leq i\leq a$. Note that |A|=a+b1|A|=a+b-1. Observe that, for every permutation $g\in G$, the set $A^{g}$ contains the block $B_{1}^{g}$, and, as $A$ intersects every block in at least one point, $A$ and $A^{g}$ have nonempty intersection. Hence, $A$ is not self-separable for $G$, and the claim follows.
From here on, the proof is divided in multiple cases. First, we prove by induction on $b$ that, if $b$ is even or if $b\geq 5$ is odd and $a\neq 3$, then $\mathbf{m}(G)=a+b-1$. The cases with $a$ or $b$ equal to $3$ are delicate, hence we need to treat them separately.
Assume that $b$ is even. Let $A$ be a subset of $a]\times[b]$ of size not exceeding a+b2a+b-2. Our base case is $b=2$. Let $\ell=|A\cap B_{1}|$, and note that $|A\cap B_{2}|\leq a$-$\ell$. Let $g\in\mathrm{Sym}(a)\mathop{\mathrm{wr}}C_{2}$ be a permutation such that the blocks $B_{1}$ and $B_{2}$ are switched. If $A$ fills one block, any choice of $g$ works. Otherwise, by $a$-transitivity of $\mathrm{Sym}(a)$, we can assure that $A\cap B_{1}\cap(A\cap B_{2})^{g}$ and $A\cap B_{2}\cap(A\cap B_{1})^{g}$ are both empty. Thus, $A$ and $A^{g}$ have trivial intersection, and $A$ is self-separable for $G$.
Now, assume that the property holds for $b$, and we claim that it also holds for $b+2$. If there does not exist a block $B\in\Sigma$, such that $|B\cap A|\geq\lceil(a+1)/2\rceil$, then, in view of Example 5, there there exists an element $g$ of the base group $\mathrm{Sym}(a)^{b}$ such that, for every $B\in\Sigma$, $|A\cap B\cap(A\cap B)^{g}|=0$. Therefore, $A$ is self-separable for $G$.
Otherwise, without loss of generality, we suppose that $|B_{1}\cap A|=\ell\geq\lceil(a+1)/2\rceil$. We aim to prove the existence of a block $B\in\Sigma$ such that $|B\cap A|\leq a$-$\ell$. Aiming for a contradiction, we assume that, for every $2\leq i\leq a$, $|B_{i}\cap A|\geq a-\ell+1$. We compute
$|A|$ $=\sum_{B_{i}\in\Sigma}|B_{i}\cap A|$
$\geq\ell+(b-1)(a-\ell+1)$
$=(b-1)(a+1)-\ell(b-2)$
$\geq(b-1)(a+1)-a(b-2)$
$=a+b-1.$
This goes against the assumption that $|A|\leq a+b-2$. Therefore, without loss of generality, we can assume that $|B_{2}\cap A|\leq a$-$\ell$. We define
$A_{0}=A\cap(B_{1}\cap B_{2})\quad\hbox{and}\quad A_{1}=A\cap\left(\bigcup_{i=3}^{b}B_{i}\right),$
and we observe that $|A_{0}|\leq a$ and that $|A_{1}|\leq a+b-\ell-2\leq a+b-4$. We can apply the inductive hypothesis on the pairs $(A_{0},\mathrm{Sym}(a)\mathop{\mathrm{wr}}C_{2})$ and $(A_{1},\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b-2))$, thus obtaining that there exists a permutation $g\in(\mathrm{Sym}(a)\mathop{\mathrm{wr}}C_{2})\times(\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b-2))\leq\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$ such that $A\cap A^{g}$ is empty. Hence, $A$ is self-separable for $G$, and the proof of this case is complete.
Assume that $b\geq 5$ is odd and that $a\neq 3$. Let $A$ be a subset of $a]\times[b]$ of size not exceeding a+b2a+b-2. By the restraint on $b$, our base case is $b=5$. We claim that $|A\cap B_{i}|>a/2$ cannot happen for three blocks at once. To fix the notation, we label this three blocks with the indices $1$, 22 and $3$. If that happens, then
$|A\cap(B_{1}\cup B_{2}\cup B_{3})|\geq 3\left\lceil\frac{a+1}{2}\right\rceil\geq a+\frac{7}{2}.$
(Observe that the last inequality does not hold if $a=3$.) This contradicts the assumption $|A|\leq a+b-2=a+3$. Hence, the number of blocks containing more than $a/2$ points is at most $2$. We need to deal with three cases based on for how many blocks this occurs. If this occurs for no blocks, we can find a permutation $g$ of the base group such that $A\cap A^{g}$ is empty, as we did for the case $b=2$. Suppose now that there exists exactly one block for which its intersection with $A$ contains more than $a/2$ points, and, without loss of generality, we assume that this block is $B_{1}$, and we let $\ell=|B_{i}\cap A|$. We aim to show that then there exists a block $B_{i}$ for which $|B_{i}\cap A|\leq b$-$\ell$. Indeed, by the way of contradiction, if, for every $2\leq i\leq 5$, $|B_{i}\cap A|\geq a-\ell+1$, then
$|A|$ $=\sum_{B_{i}\in\Sigma}|B_{i}\cap A|$
$\geq\ell+4(a-\ell+1)$
$=4a-3\ell+4$
$\geq a+4.$
This goes against the cardinality of $A$. Without loss of generality, $|B_{2}\cap A|\leq a$-$\ell$. Hence, reasoning as in the previous base case, we see that there exists an element $g$ such that $g$ stabilizes the blocks $B_{3}$, $B_{4}$ and $B_{5}$, it swaps the blocks $B_{1}$ and $B_{2}$ and $A\cap A^{g}$ is empty. A similar argument can be carried out to deal with the remaining case. Henceforth, we have proved that every subset $A$ containing less that $a+4$ points is self-separable for $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(5)$ (and $a\neq 3$).
The inductive step for $b$ odd (and $a\neq 3$) is verbatim the one we used for $b$ even.
Assume that $b=3$ and $a$ is odd. We first show that $\mathbf{m}(G)\leq a+b-2=a+1$. Indeed, we choose $A$ so that
$|B_{1}\cap A|=|B_{2}\cap A|=\frac{a+1}{2}.$
Then, for every $g\in G$, $A^{g}$ and $A$ will intersect in at least one point in $B_{1}$ or $B_{2}$. Hence $A$ is not self-separable. Our objective is to show that every set $A$ whose cardinality is at most $a$ is self-separable for $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(3)$.
We claim that, up to permuting the indices of the blocks, if $|B_{1}\cap A|=\ell\geq(a+1)/2$, then $|B_{2}\cap A|\leq a$-$\ell$, and $|B_{3}\cap A|\leq(a+1)/2$. Aiming for a contradiction, assume that either $|B_{2}\cap A|\leq a$-$\ell$ or $|B_{3}\cap A|\leq(a+1)/2$ fails. If the former happens, we see that
$|A|\geq|B_{1}\cap A|+|B_{2}\cap A|>\ell+a-\ell=a.$
Meanwhile, if the latter holds, then
$|A|\geq|B_{1}\cap A|+|B_{3}\cap A|\geq 2\frac{a+1}{2}\geq a+1.$
In both cases, we find a contradiction with $|A|\leq a$. Using our usual arguments, our claim implies that there exists a permutation $g\in\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(3)$ such that $g$ swaps the blocks $B_{1}$ and $B_{2}$, it fixes $B_{3}$ and $A\cap A^{g}$ is empty. Therefore, we established the theorem in this setting.
Assume that $b=3$ and $a$ is even. Let $A$ be a subset of cardinality at most $a+1$. We can deal with the case with every block containing less that $a/2$ points as usual. Hence, without loss of generality, we suppose that $|B_{1}\cap A|=\ell\geq a/2+1$. It cannot be that, for every $i\in{2,3}$, $|B_{2}\cap A|\geq a-\ell+1$, otherwise
$|A|\geq\sum_{B_{i}\in\Sigma}|A\cap B_{i}|=\ell+2(a-\ell+1)\geq a+2,$
[…]
Assume that $b$ is odd and $a=3$. Observe that, in this setting, there exists a set of size $b+1$ which is not self-separable. Indeed, let $A$ be a subset of $a]\times[b]$ such that, for every $i\leq\lceil(b+1)/2\rceil$, $|A\cap B_{i}|=2$, and $A\cap B_{i}$ is empty otherwise. For every permutation $g\in\mathrm{Sym}(3)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$, there exists an index $i$ such that
$|(A\cap B_{i})\cap(A\cap B_{i})^{g}|\geq 1.$
Hence $\mathbf{m}(G)\leq b+1$.
Finally, we argue by induction on $b$. The base case $b=3$ was already taken care of before, while the induction argument is the same we used already two times in this proof.
This last case exhaust the possibilities for the pairs $(a,b)$, and hence it completes the proof of the first part of the statement of I.
If $G$ is an imprimitive permutation group, and $\Sigma$ is a system of blocks such that $|\Sigma|=b$ and each block contains $a$ point, then $G$ embeds in $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$. The second part of the statement then follows combining the first one with Lemma 12. ∎
7. Sharpness results for D
In this section, we characterize the permutation groups of degree $n$ with $\mathbf{m}(G)=\lceil(n+1)/2\rceil$ – that is, we characterize the permutation groups that meet the upper bound from D.
We divide the discussion into two fundamentally distinct cases: primitive groups (see Section 7.1) and imprimitive groups (see Section 7.2). Namely, we prove the following characterizations.
Lemma 21.
Let $G$ be a primitive permutation group of degree $n$. Then $\mathbf{m}(G)=\lceil(n+1)/2\rceil$ if, and only if, one of the following possibilities holds:
  • $n\geq 2$ and $G\cong\mathrm{Sym}(n)$ in natural action;
  • $n\geq 3$ and $G\cong\mathrm{Alt}(n)$ in natural action;
  • $n=5$ and $G\cong{\rm C}_{5}$, $\rm Dih(5)$}, ${\rm AGL}_{1}(5)$;
  • $n=6$ and $G\cong\mathrm{Sym}(5)$ in its action on $6$ points;
  • $n=7$ and $G\cong{\rm Dih}(7)$, ${\rm AGL}_{1}(7)$;
  • $n=8$ and $G\cong{\rm AGL}_{1}(8)$, ${\rm A\Gamma L}_{1}(8),{\rm ASL}_{3}(2)$, ${\rm PSL}_{2}(7)$, ${\rm PGL}_{2}(7)$;
  • $n=9$ and $G\cong{\rm AGL}_{1}(9)$, ${\rm PSU}_{3}(2)\cong M_{9}$, ${\rm A\Gamma L}_{1}(9)$, ${\rm ASL}_{2}(3)$, ${\rm AGL}_{2}(3)$, ${\rm PSL}_{2}(8)$, ${\rm P\Gamma L}_{2}(8)$;
  • $n=10$ and $G\cong{\rm PGL}_{2}(9)$, ${\rm P\Gamma L}_{2}(9)$;
  • $n=12$ and $G\cong{\rm PGL}_{2}(11)$, ${\rm P\Gamma L}_{2}(11)$, $M_{11}$, $M_{12}$;
  • $n=14$ and $G\cong{\rm PGL}_{2}(13)$;
  • $n=16$ and $G\cong{\rm AGL}_{4}(2)$;
  • $n=24$ and $G\cong M_{24}$.
Lemma 22.
Let $G$ be an imprimitive permutation group of degree $n$ such that $\mathbf{m}(G)=\left\lceil(n+1)/2\right\rceil$. Then, $n$ is even, and $G$ defines a block system $\Sigma$ with either $n/2$ blocks of size $2$ or two blocks of size $n/2$. In particular, one of the following possibilities holds:
  1. $(a)$
    $n\leq 24$ and $G$ is permutationally isomorphic to a group TransitiveGroup(n,k)(n,k) from the library of transitive permutation groups in Magma [9] (based on [15, 23, 24]), where the pairs $(n,k)$ are recorded in Table 2;
  2. $(b)$
    $n\geq 16$ and the permutation group that $G$ induces on the block system $\Sigma$ is isomorphic to $\mathrm{Alt}(n/2)$ or $\mathrm{Sym}(n/2)$ in natural action (see Lemma 27 for the list of groups);
  3. $(c)$
    $n\geq 16$ and the permutation group that $G$ induces on a given block of $\Sigma$ isomorphic to $\mathrm{Alt}(n/2)$ or $\mathrm{Sym}(n/2)$ in natural action (see Lemma 28 and the previous discussion for the list of groups).
Degree $n$ # of groups Database numbers $k$
44 33 11, 22, 33
66 1919 22, 33, 55, 88, 99, $10$, $11$, $13$, $17$, $18$, 1919, 2121, $22$, $26$, $28$, $29$, $30$, $31$, $35$
88 $28$ 33, 99, $13$, $14$, $17$, $18$, 1919, 2121, $22$, $24$, $26$, $28$, $29$, $30$, $31$, $32$, $33$, $34$, $35$, $38$, $39$, $40$, $41$, $42$, $44$, $45$, $46$, $47$
$10$ $17$ $11$, $12$, $17$, 1919, 2020, $22$, $25$, $27$, $29$, $33$, $36$, $38$, $39$, $40$, $41$, $42$, $43$
$12$ $22$ $123$, $180$, $181$, $182$, $183$, $219$, $220$, $256$, $257$, $270$, $277$, $278$, $279$, $285$, $286$, $287$, $288$, $293$, $296$, $297$, $298$, $299$
$14$ $10$ $46$, $47$, $49$, $54$, $56$, $57$, $58$, $59$, $60$, $61$
$16$ $23$ $1078$, $1502$, $1506$, $1693$, $1798$, $1799$, $1801$, $1802$, $1803$, $1804$, $1805$, $1842$, $1843$, $1844$, $1860$, $1861$, $1878$, $1882$, $1883$, $1902$, $1903$, $1916$, $1940$
$18$ 66 $855$, $897$, $914$, $937$, $938$, $952$
$24$ 44 $18439$, $24652$, $24732$, $24924$
Table 2. List of imprimitive permutation groups of degree at most $24$ (with degree $14$ added) that meet the bound of D.
7.1. Primitive groups
We start our analysis with two general observations, that will greatly reduce the possibilities for primitive permutation group meeting the bound from D.
Lemma 23.
Let $G$ be a permutation group on $\Omega$ of degree $n$, let $c(g)$ be the number of cycles in the cyclic decomposition of $g\in G$, let $\ell(g)$ be the length of the shortest odd cycle in that composition (with $\ell(g)=0$ if $g$ contains no odd cycles), and for a non-negative integer $i$, let $\mathcal{O}_{i}$ be the set of elements $g$ in $G$ whose cycle decomposition contains at most $i$ odd cycles. Further, let $r(G)$ be the number of orbits of $G$ in its induced action on $\Omega\choose\lfloor n/2\rfloor$}. Suppose that $\mathbf{m}(G)=\lceil(n+1)/2\rceil$. Then, if $n$ is even,
$r(G)|G|=\sum_{A\in{\Omega\choose n/2}}|G_{A}|=\sum_{g\in\mathcal{O}_{0}}2^{c(g)},$
and, if $n$ is odd,
$r(G)|G|=\sum_{A\in{\Omega\choose\lfloor n/2\rfloor}}|G_{A}|\leq\sum_{g\in\mathcal{O}_{1}}2^{c(g)-1}\ell(g)\leq\left\lceil\frac{n}{2}\right\rceil\sum_{A\in{\Omega\choose\lfloor n/2\rfloor}}|G_{A}|=\left\lceil\frac{n}{2}\right\rceil r(G)|G|.$
Proof.
Let us start by observing that the equality
$r(G)|G|=\sum_{A\in{\Omega\choose\lfloor n/2\rfloor}}|G_{A}|$
appearing in both formulas, for $n$ odd and for $n$ even, is just a restatement of the Cauchy-Frobenius Lemma (see [18, Theorem 1.7A]).
In what follows, we shall assume without loss of generality that $\Omega=[n]$. By double counting the set
$\left\{(A,g)\in{[n]\choose\lfloor n/2\rfloor}\times G\>\middle|\>A\cap A^{g}=\emptyset\right\},$
we see that
(F) $\sum_{A\in{[n]\choose\lfloor n/2\rfloor}}\left|\left\{g\in G\mid A\cap A^{g}=\emptyset\right\}\right|=\sum_{g\in G}\left|\left\{A\in{[n]\choose\lfloor n/2\rfloor}\>\middle|\>A\cap A^{g}=\emptyset\right\}\right|.$
First, we focus on the case when $n$ is even. Since we assume that every $(n/2)$-subset AA is self-separable, and there is a unique image of $A$ that is disjoint from it, it follows that $\left\{g\in G\mid A\cap A^{g}=\emptyset\right\}=G_{A}h$, where $h\in G$ is any element such that $A\cap A^{h}=\emptyset$. Hence,
$\sum_{A\in{[n]\choose\lfloor n/2\rfloor}}\left|\left\{g\in G\mid A\cap A^{g}=\emptyset\right\}\right|=\sum_{A\in{[n]\choose\lfloor n/2\rfloor}}|G_{A}|.$
Let us now count how many $(n/2)$-subsets can be separated from themselves by a fixed $g\in G$. Suppose $g$ is written as a product of $c(g)$ disjoint cycles, and let $A$ be such a set. Observe that $A$ must meet every cycle of the decomposition in precisely half of the points of that cycle. In particular, every cycle has even length. Furthermore, if $\pi=(\pi_{1},\pi_{2},\dots,\pi_{2\ell})$ is one of the cycles of $g$, AA must contain either all the $\pi_{i}$ whose indices are odd or all those whose indices are even. In particular, we can choose in two ways which points of each of the $c(g)$ cycles to include in $A$. Hence, there exist $2^{c(g)}$ sets $A$ such that $|A|=n/2$ and $A\cap A^{g}$ is empty. This gives us
$\sum_{g\in G}\left|\left\{A\in{[n]\choose n/2}\>\middle|\>A\cap A^{g}=\emptyset\right\}\right|=\sum_{g\in\mathcal{O}_{0}}2^{c(g)},$
proving the lemma in the case when $n$ is even.
Suppose now that $n$ is odd. We start by considering the left-hand side of Equation F. Fix a set $A\in{[n]\choose\lfloor n/2\rfloor}$. Let ${\mathcal{B}}_{A}=\{A^{g}\mid g\in G,A\cap A^{g}=\emptyset\}$, and for each $B\in{\mathcal{B}}_{A}$ choose an element $h_{B}\in G$ mapping $A$ to $B$. Let $T=\{h_{B}:B\in{\mathcal{B}}_{A}\}$. Clearly, $|T|=|{\mathcal{B}}_{A}|\leq\lceil n/2\rceil$. Similarly as in the even case, we see that for each $B\in{\mathcal{B}}_{A}$, the set of elements $g\in G$ mapping $A$ to $B$ equals $G_{A}h_{B}$. Now observe that
$\left|\left\{g\in G\mid A\cap A^{g}=\emptyset\right\}\right|=\left|\left\{g\in G\mid A^{g}\in{\mathcal{B}}_{A}\right\}\right|=\sum_{B\in{\mathcal{B}}_{A}}\left|\left\{g\in G\mid A^{g}=B\right\}\right|=|{\mathcal{B}}_{A}|\,|G_{A}|.$
Therefore the left-hand side of Equation F is equal to
$\sum_{A\in{[n]\choose\lfloor n/2\rfloor}}|{\mathcal{B}}_{A}||G_{A}|,$
and since $1\leq|{\mathcal{B}_{A}}|\leq\lceil n/2\rceil$, it is bounded below and above by the quantities
[…]
respectively.
It remains to show that the right-hand side of Equation F is equal to $\sum_{g\in\mathcal{O}_{1}}2^{c(g)-1}\ell(g)$. Again, for a given permutation $g\in G$, we count all the possible $\lfloor n/2\rfloor$-subsets $A$ such that $A$ and $A^{g}$ are disjoint. The situation is slightly more complex in this case. In the same way as before, we observe that, if $g$ contains more than one odd cycle, such an $A$ does not exist. Hence, we assume that $g$ contains precisely one odd cycle of length $\ell(g)$, and c(g)1c(g)-1 even cycles. Note that $A$ intersects the odd cycle in precisely $\lfloor\ell(g)/2\rfloor$ points. Moreover, no two consecutive points of the odd cycle are contained in $A$, implying that precisely two consecutive points are excluded from $A$. Once these two consecutive points are determined, the remaining points of the cycle contained in $A$ are uniquely determined (by the requirement that they need to alternate between being contained and not in $A$). Since there are $\ell(g)$ choice for the two consecutive points excluded from $A$, and repeating the same reasoning as before for the points belonging to an even cycle, we obtain
$\sum_{g\in\mathcal{O}_{1}}\left|\left\{A\in{[n]\choose\lfloor n/2\rfloor}\>\middle|\>A\cap A^{g}=\emptyset\right\}\right|=\sum_{g\in\mathcal{O}_{1}}2^{c(g)-1}\ell(g).$
This concludes the proof. ∎
Given two functions $f,g:\mathbb{N}\to\mathbb{R}^{+}$, we write $f\sim g$ if the limit, as $n$ goes to infinity, of the quotient $f(n)/g(n)$ is $1$.
Lemma 24.
Suppose that $G$ is a permutation group of degree $n$ such that $\mathbf{m}(G)=\lceil(n+1)/2\rceil$. Then, if $n$ is even,
$|G|\geq\frac{1}{2^{n/2}}{n\choose n/2}\sim\frac{1}{\sqrt{2\pi}}\cdot 2^{\frac{1}{2}(n-\log_{2}n)},$
while, if $n$ is odd,
$|G|\geq\frac{1}{2^{\lfloor n/2\rfloor}n}{n\choose\lfloor n/2\rfloor}\sim\sqrt{\frac{2}{\pi}}\cdot 2^{\frac{1}{2}(n-3\log_{2}n)}.$
Proof.
These bounds are obtained by disregarding the exclusion of the elements from $\mathcal{O}_{i}$, by substituting in the formulae from Lemma 23 the lower bound $1$ for $|G_{A}|$ and the upper bound $2^{n/2}$ for $2^{c(g)}$ and $n$ for $\ell(g)$, and by applying Stirling’s formula to derive the asymptotic expansion. ∎
We are ready to characterize which primitive permutation groups meet the putative upper bound.
Proof of Lemma 21.
By [32, Theorem 1.1], a primitive group whose degree exceeds $24$ is either a subgroup of $\mathrm{Sym}(a)\mathop{\mathrm{wr}}\mathrm{Sym}(b)$ with socle isomorphic to $\mathrm{Alt}(a)^{b}$, where the action of $\mathrm{Sym}(a)$ is on $k$-subsets (for some $1\leq k\leq n/2)$, the action of $\mathrm{Sym}(b)$ in on $b$ points, and the wreath product is endowed with the product action, or
(G) $|G|\leq n^{1+\log_{2}n}.$
Using Wolfram Mathematica, we compute that the function appearing in Lemma 24 is greater that the bound from Equation G whenever $n$ is an even integer greater or equal to $116$, or $n$ is an odd integer greater or equal to $147$.
We have two cases to consider. If the degree of the permutation group $G$ exceeds $116$ or $147$, depending on the parity of its degree, by L, for every $a\geq 4$ and $b\geq 2$,
$\mathbf{m}(\mathrm{Sym}(a)\mathop{\mathrm{wr}}Sym(b))\leq\left\lceil\frac{a+1}{2}\right\rceil^{b}<\left\lceil\frac{a^{b}+1}{2}\right\rceil.$
In particular, if the degree exceeds $16$, we can assume that $G$ is of almost simple type. Moreover, Lemma 29 implies that $G$ is permutationally isomorphic to either $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$ endowed with their actions on $n$ points.
Let us now deal with the groups of small degree, where we can use a computational approach. We started with a list of all primitive permutation groups of degree at most $146$ in the database of primitive groups in Magma, where we excluded the symmetric groups and the alternating groups in their natural actions from consideration. (This database is based on [17, 41, 42, 44].) Each of these primitive groups of even degree was tested against the inequality
${n\choose\frac{n}{2}}\leq\sum_{g\in\mathcal{O}_{0}}2^{c(g)},$
while those of odd degree against
${n\choose\frac{n-1}{2}}\leq\sum_{g\in\mathcal{O}_{1}}2^{c(g)-1}\ell(g).$
Both inequalities are obtained from Lemma 23, by substituting to each $|G_{A}|$ a $1$ in the sum on the right hand side of the expressions.
Our search list get significantly shortened: the possible groups have degree at most $32$, with the only candidate of degree $32$ being ${\rm AGL}(5,2)$. We use a different approach for it. By randomly sampling subsets of size at most $16$ of the domain $\mathbb{F}_{2}^{5}$ and then testing them for existence of a group element that maps them to their complement, we found a number of sets of size $15$ that are not self-separable, thus showing that $\mathbf{m}({\rm AGL}(5,2))\leq 15$ (and we believe that $15$ is the exact value). The remaining groups all have degree $24$ or less, with the only groups of order $24$ being the Mathieu group $M_{24}$. These groups were dealt with an exhaustive search of $\lfloor(n+1)/2\rfloor$-subsets which are not self-separable. The groups where such a set was found were excluded form the list, and those that survived are precisely the ones listed in Lemma 21. Our code for these operations can be found in [5]. This concludes the proof. ∎
7.2. Imprimitive groups
We start this section with the following lemma, of independent interest, which gives a necessary condition for a permutation group to be $k$-homogeneus.
Lemma 25.
Let $G$ be a transitive permutation group acting on a set $\Omega$ of size $n$, and let $k\leq\lceil n/2\rceil-1$. Suppose that, for every two disjoint subsets $A,B\subseteq\Omega$ of size $k$, there exists $g\in G$ mapping $A$ to $B$. Then, $G$ is $k$-homogeneous.
Proof.
Aiming for a contradiction, suppose that the claim of the lemma is false. Then the set
$\mathcal{E}=\{(A,B)\mid A,B\subseteq\Omega,|A|=|B|=k,A^{g}\neq B\text{ for every }g\in G\}$
is nonempty. Among all $(A,B)\in\mathcal{E}$ we choose one which minimises $|A\cap B|$. Observe that by our assumptions, the size $t$ of the intersection $A\cap B$ satisfies $1\leq t\leq k-1$. Thus,
$|\Omega\setminus(A\cup B)|=n-2k+t\geq t+1.$
This allows us to choose a subset $C\subseteq\Omega\setminus(A\cup B)$ with $|C|=t$. Now, let
$A^{\prime}=(B\setminus A)\cup C$
and observe that $A\cap A^{\prime}$ is empty. Hence, by the assumptions, there exists $g\in G$ such that $A^{g}=A^{\prime}$.
Observe that the sets $A\setminus B$ and $\Omega\setminus(A\cup B\cup C)$ are disjoint and of respective cardinalities ktk-t and n2kn-2k. Their union is therefore of size nktn-k-t. If $n-k-t\geq k$, then let $A^{\prime\prime}$ be an arbitrary $k$-subset of this union. Note that such a set $A^{\prime\prime}$ is disjoint from $A^{\prime}$, implying that there exist $g^{\prime}\in G$ such that $A^{\prime g^{\prime}}=A^{\prime\prime}$. Observe also that $B\cap A^{\prime\prime}=\emptyset$, implying that there exists $h\in G$ such that $A^{\prime\prime h}=B$. But then $A^{gg^{\prime}h}=B$, contradicting the fact that $(A,B)\in\mathcal{E}$. This contradiction implies that nkt<kn-k-t<k, or equivalently, 2k+tn=k(nkt)>02k+t-n=k-(n-k-t)>0. Since $2k\leq n-1$, it follows that $2k+t-n\leq t-1$, allowing us to choose a (2kn+t)(2k-n+t)-subset $D$ of $A\cap B$. Now let
$A^{\prime\prime}=(A\setminus B)\cup(\Omega\setminus(A\cup B\cup C))\cup D,$
and observe that $|A^{\prime\prime}|=k$ and $A^{\prime\prime}\cap A^{\prime}=\emptyset$. Similarly as above, there exist $g^{\prime}\in G$ such that $A^{\prime g^{\prime}}=A^{\prime\prime}$. Observe also that $|B\cap A^{\prime\prime}|=|D|<t$, implying that there exists $h\in G$ such that $A^{\prime\prime h}=B$, which contradicts the fact that $(A,B)\in\mathcal{E}$. ∎
Remark 26.
Note that the upper bound on $k$ in the assumptions of the above lemma cannot be improved to $\lceil n/2\rceil$ since that would imply that every transitive permutation group $G$ of even degree $n$ and with $\mathbf{m}(G)=n/2$ is $k$-homogeneous. However, as we show later in this section, this is far from true.
From the remainder of this section, we prove Lemma Lemma 22. Let $G$ be an imprimitive permutation group of degree $m$ with $\mathbf{m}(G)=\lceil(n+1)/2\rceil.$ We observe that I implies that $G$ admits a block system consisting either of two blocks of size $m/2$, or of $m/2$ blocks of size 2. Since $m$ is even, we write $m=2n$ throughout this proof.
Suppose that $G$ is imprimitive with blocks of size $2$. (We also suppose that the degree is at least $6$, and we will deal with the smaller degrees computationally.) Let $\Sigma$ be the corresponding system of imprimitivity, consisting of $n\geq 3$ blocks of size $2$. Note that, for every subset $A$ of the domain, $\Sigma$ can be partitioned in three sets, $\Sigma=\Sigma_{0}\cup\Sigma_{1}\cup\Sigma_{2}$, such that every block in $\Sigma_{i}$ intersects $A$ in $i$ points and since every subset $A\subseteq\Omega$ of size $n$ is self-separable, $|\Sigma_{0}|=|\Sigma_{2}|=\lceil n/3\rceil$. Since $A$ is self-separable, there exists a permutation $g\in G$ which maps $A$ to its complement, and thus maps $\Sigma_{0}$ bijectively to $\Sigma_{2}$, and vice versa. In particular, $H:=G^{\Sigma}$, the permutation group $G$ induced on the block system $\Sigma$, contains an element that maps $\Sigma_{0}$ in $\Sigma_{2}$. If we allow $A$ to range through all the subsets of cardinality $n$ and satisfying $|\Sigma_{0}|=|\Sigma_{2}|=\lceil n/3\rceil$, we conclude that $G^{\Sigma}$ is $\lceil n/3\rceil$-homogeneous by Lemma 25. The list of permutation groups with this property is quite short. We recall that [18, Theorem 9.4B] states that, if $H$ is a finite and transitive $k$-homogeneous permutation group, then $H$ is either $k$-transitive or it belongs to a well-understood list of exceptions. Note that all $k$-homogeneous groups are also $2$-transitive, so combining this with the classification of $2$-transitive permutation groups (see [13, Section 5]), if $H$ is $\lceil n/3\rceil$-homogeneous and primitive, the list of possibilities for $H$ is as follows:
  1. $(a)$
    $\mathrm{Alt}(n)\leq H\leq\mathrm{Sym}(n)$ in their natural action on $n$ points;
  2. $(b)$
    $\mathrm{ASL}_{d}(q)\leq H\leq\mathrm{A}\Gamma\mathrm{L}_{d}(q)$ in their natural action on $q^{d}$ points, where
    $(d,q)\in\{(1,4),(1,5),(1,8),(2,2),(3,2)\};$
  3. $(c)$
    $\mathrm{PSL}_{d}(q)\leq H\leq\mathrm{P}\Gamma\mathrm{L}_{2}(q)$ in their natural action on $\mathrm{PG}_{1}(q)$, where $q\in\{4,5,7,8\}$;
  4. $(d)$
    $H$ is isomorphic to either $M_{11}$ or $M_{12}$ in their natural actions on $11$ or $12$ points, respectively.
Hence, to prove our lemma we need to consider these cases in turn. We will handle cases  $(b),$  $(c)$ and  $(d)$ with $H\cong M_{11}$ by computation at the end of the proof. Now we will provide some useful structural descriptions for cases  $(a)$ and  $(d)$ with $M_{12}$. In case $(a)$, we will use these explicit description to determine which groups meet the upper bound in Lemma 27. For  $(d)$ with $M_{12},$ these will allow us to restrict the search space of the algorithm to a manageable size, so we can handle this but computation at the end of the proof as well.
Let $H$ be a permutation group acting on a finite set $\Omega$, and let $\mathbb{F}_{2}\Omega$ be the permutation module of $H$ over the field $\mathbb{F}_{2}$ with two elements. Note that the wreath product $C_{2}\mathop{\mathrm{wr}}H$ acts imprimitively on $\mathbb{F}_{2}\Omega$. Every imprimitive permutation group $G$ such that a block system $\Sigma$ for $G$ consists of blocks of size $2$, and the permutation group that $G$ induces on $\Sigma$ is isomorphic to $H$, embeds into $C_{2}\mathop{\mathrm{wr}}H$. Moreover, the kernel $K$ of the action of $G$ on $\Sigma$ is an $\mathbb{F}_{2}H$-submodule of $\mathbb{F}_{2}\Omega$, otherwise $K$ would not be normal in $G$. The possibilities for $K$ are, as a consequence, quite limited. Indeed, in both cases, [35] states that $K$ is either trivial, the one-dimensional vector space $D$ spanned by the all-$1$ vector, the (n1)(n-1)-dimensional vector space $A$ containing all the vector whose entries sum equals to zero, and the whole module $\mathbb{F}_{2}\Omega$. We now divide our discussion according to the group $H$.
Suppose that $H=M_{12}$.
  • If $K$ is trivial, $G$ is isomorphic to $H$ as abstract groups. On the other hand, as the blocks of $\Sigma$ have cardinality $2$ and as the stabiliser in $H$ is the stabilizer of a block in $G$, the stabiliser in $G$ has index $2$ in the stabiliser in $H$. As the stabilizer in $H$ is isomorphic to $M_{11}$, which is simple, it contains no subgroups of index $2$. Hence no such permutation group $G$ exists.
  • If $K$ is isomorphic to $D$, $G$ is isomorphic, as abstract groups, to either $C_{2}\times H$ or to the central perfect extension of $H$ by $C_{2}$. In the former case, arguing as before, we found a single possible stabilizer in $G$, which isomorphic to $M_{11}$. In the latter, we observe that, since the Schur multiplier of $M_{11}$ is trivial, the stabilizer of a block of $\Sigma$ is isomorphic to $C_{2}\times M_{11}$. As a consequence, $M_{11}$ is the unique (up to conjugation) subgroup of the central perfect extension, which enjoys the property of being a stabiliser in $G$.
  • If $K$ is isomorphic to $A$, we claim that $G$ is isomorphic, as abstract groups, to the split extension $A\rtimes H$. In fact, if this were not the case, $A\rtimes H$ would have an index $2$ subgroup, which goes against the simplicity of $H$ and the possibilities for $K$. The stabilizer in $G$ is unique (up to conjugation): this follows from the uniqueness of the subgroup $\mathbb{F}_{2}\Omega\rtimes M_{11}$ of index $12$ in the wreath product $\mathbb{F}_{2}\Omega\rtimes M_{12}$, and that, as the action on the blocks is transitive, $H$ acts transitively on the subgroups of index $2$ of $\mathbb{F}_{2}\Omega$. In particular, the stabiliser can be characterized as the intersection of $G$ with the stabilizer of the whole wreath product endowed with the imprimitive action.
  • If $K$ is isomorphic to $\mathbb{F}_{2}\Omega$, the same reasoning as the previous case shows that $G$ coincides with the whole wreath product endowed with the imprimitive action.
Suppose that $H=\mathrm{Alt}(n)$, with $n\geq 8$. (The cases with $n\leq 7$ will be dealt with computationally.) The same reasoning as before can be applied if $K\in\{1,A,\mathbb{F}_{2}\Omega\}$.
  • If $K$ is trivial, we find no possibilities for $G$.
  • If $K$ is isomorphic to $A$, $G$ is the split extension $A\rtimes H$, and the stabilizer is obtained by intersecting $G$ with the stabilizer of the whole wreath product endowed with the imprimitive action.
  • If $K$ is isomorphic to $\mathbb{F}_{2}\Omega$, $G$ is with the whole wreath product endowed with the imprimitive action.
The case $K=D$ is slightly different, as our hypothesis on the degree prescribes that the Schur multipliers of both $\mathrm{Alt}(n)$ and its stabiliser $\mathrm{Alt}(n-1)$ are cyclic of order $2$.
  • If $K$ is isomorphic to $D$, $G$ is isomorphic, as abstract groups, to either $C_{2}\times H$ or to the central perfect extension of $H$ by $C_{2}$. In the former case, as before, $\mathrm{Alt}(n-1)$ is the single possibility for the stabiliser in $G$ (up to conjugation). In the latter, the stabiliser of a block of $\Sigma$ is isomorphic to the central perfect extension of $\mathrm{Alt}(n-1)$ by $C_{2}$. As $\mathrm{Alt}(n-1)$ is simple, this extension does not have any index two subgroup. Hence, if $G$ is the central perfect extension of $H$ by $C_{2}$, $G$ does not admit an imprimitive action of degree $2n$.
Finally, we suppose that $H=\mathrm{Sym}(n)$, with $n\geq 8$. The fact that $H$ contains $\mathrm{Alt}(n)$ as a subgroup of index $2$ makes the treatment of this case more convoluted.
  • If $K$ is trivial, $G$ is isomorphic to $H$ as abstract groups, and the stabilizer in $G$ is the unique (up to conjugation) subgroup of index $2n$ in $\mathrm{Sym}(n)$, that is, $\mathrm{Alt}(n-1)$.
  • If $K$ is isomorphic to $D$, $G$ is isomorphic, as abstract groups, to either $C_{2}\times H$ or to the central perfect extension of $H$ by $C_{2}$. In the former case, as the stabiliser of a block of $\Sigma$ is $C_{2}\times\mathrm{Sym}(n-1)$, there are two subgroups of index $2$ of $C_{2}\times\mathrm{Sym}(n-1)$: it is isomorphic to either $\mathrm{Sym}(n-1)$ or to $\mathrm{Alt}(n-1)\times C_{2}$. Only the first one can be the stabiliser in $G$, because $\mathrm{Alt}(n-1)\times C_{2}$ is not core-free. In the latter, we observe that the stabilizer of a block of $\Sigma$ is isomorphic to the central perfect extension of $\mathrm{Sym}(n-1)$ by $C_{2}$. The unique subgroup of index $2$ of the stabiliser of a block is the central perfect extension of $\mathrm{Alt}(n-1)$ by $C_{2}$, which is not core-free. Hence, $G$ cannot be endowed with an imprimitive action on $2n$ points.
  • If $K$ is isomorphic to $A$, we observe that $A\rtimes\mathrm{Alt}(n)$ has index $4$ as a subgroup of $C_{2}\mathop{\mathrm{wr}}\mathrm{Sym}(n)$, and their quotient is isomorphic to $C_{2}\times C_{2}$. Hence, there are three possibilities for $G$ as an abstract group. If $G$ is isomorphic to the split extension $A\rtimes H$, as we did for $M_{12}$ and $\mathrm{Alt}(n)$, we conclude that the stabilizer in $G$ is isomorphic to the intersection of $G$ with the stabiliser in $C_{2}\mathop{\mathrm{wr}}\mathrm{Sym}(n)$. If $G$ is isomorphic to $C_{2}\mathop{\mathrm{wr}}\mathrm{Alt}(n)$, we note that $G$ induces $\mathrm{Alt}(n)$ as permutation group on $\Sigma$, against our hypothesis. Last, if $s$ denotes an element of the wreath product outside of $A$, the remaining possibility for $G$ is generated as
    $G=\left\langle A,\mathrm{Alt}(n),sg\mid g\in\mathrm{Sym}(n)\setminus\mathrm{Alt}(n)\right\rangle,$
    and its stabilizer is obtained by intersecting $G$ with the whole wreath product.
  • If $K$ is isomorphic to $\mathbb{F}_{2}\Omega$, $G$ coincides with the whole wreath product endowed with the imprimitive action.
We are ready to determine for which groups arising in $(a)$ the upper bound from D is attained.
Lemma 27.
Let $n\geq 8$ be a positive integer, and let $G$ be an imprimitive group with block system $\Sigma$ that consists of $n$ blocks of size $2$. Suppose that the action of $G$ on $\Sigma$ is permutationally isomorphic to $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$. Then $\mathbf{m}(G)=n+1$ if, and only if, $G$ is one of the following:
  1. $(i)$
    $\mathrm{Sym}(n)$ with stabiliser $\mathrm{Alt}(n-1)$ acting on $2n$ points;
  2. $(ii)$
    $C_{2}\times\mathrm{Alt}(n)$;
  3. $(iii)$
    $C_{2}\times\mathrm{Sym}(n)$;
  4. $(iv)$
    $A\rtimes\mathrm{Alt}(n)$ for even $n$;
  5. $(v)$
    $A\rtimes\mathrm{Sym}(n)$ for even $n$;
  6. $(vi)$
    $\langle A,\mathrm{Alt}(n),sg\mid g\in\mathrm{Sym}(n)\setminus\mathrm{Alt}(n)\rangle$, where $A$, $\mathrm{Alt}(n)$ and $s$ are as before;
  7. $(vii)$
    $C_{2}\mathop{\mathrm{wr}}\mathrm{Alt}(n)$ endowed with the imprimitive action;
  8. $(viii)$
    $C_{2}\mathop{\mathrm{wr}}\mathrm{Sym}(n)$ endowed with the imprimitive action;
Proof.
We start with the only subgroups of $C_{2}\mathop{\mathrm{wr}}H$ (where $H$ is $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$) that induce $H$ on the blocks and for which we show that they do not meet the upper bound.
We analyze in turn the possible groups we have described in the discussion preceding this lemma. In what follows, $X$ denotes a $k$-subset of the domain, with $k\leq n$, and $\Sigma_{0},\Sigma_{1}$ and $\Sigma_{2}$ denote the sets of blocks that contain 0, 11 and $2$ points from $X$, respectively.
Let us start with those cases where $K$, the kernel of the action on the block system $\Sigma$, is trivial. The only possibility is that $G=\mathrm{Sym}(n)$, and its stabiliser is $\mathrm{Alt}(n-1)$. In particular, $\mathrm{Alt}(n)\leq G$ acts intransitively defining two orbits of size $n$, while every odd permutation, along with its action on the $n$ blocks, swaps the two elements within each block. By $n$-transitivity of the action of $G$ on the blocks, there exists an odd permutation $g$ that exchanges $\Sigma_{0}$ and $\Sigma_{2}$. (Since $n\geq 4$, if $g^{\prime}$ is an even permutation with the property, then $g$ is an odd permutation obtained by multiplying $g^{\prime}$ by a transposition that exchanges two blocks in a $\Sigma_{i}$, for some $i\in\{0,1,2\}$). By our description of $G$, it is immediate to verify that $X\cap X^{g}$ is empty. This completes the proof of $(i)$. Note that $(vi)$ contains a subgroup permutationally isomorphic to $G$, and hence it enjoys the required property.
Suppose that $K=D$ is the diagonal subspace. If $H$, the group that $G$ induces on $\Sigma$, is alternating, we can determine $g$ such that $X\cap X^{g}$ is empty using similar ideas as before. By (n2)(n-2)-transitivity of $\mathrm{Alt}(n)$, there exists a permutation $h$ that swaps the sets $\Sigma_{0}$ and $\Sigma_{2}$. If we compose $h$ with $s$, the nontrivial element of $D$ that swaps the elements within each blocks, we immediately see that $X\cap X^{hs}$ is empty. Thus, $g=hs$, and $(ii)$ is obtained. Observe that $(iii)$, $(vii)$ and $(viii)$ follows immediately as $\mathrm{Alt}(n)\times C_{2}$ is a subgroup of the permutation groups into consideration. Note that we have exhausted all the cases with $K\in\{D,\mathbb{F}_{2}\Omega\}$.
Last, we suppose that $K=A$. We need to consider two cases according to $n$ being even or odd. Observe that $C_{2}=D\leq A$ if, and only if, $n$ is even. Therefore, if $n$ is even $\mathrm{Alt}(n)\times C_{2}$ is a subgroup of the groups in question: this takes care of $(iv)$ and $(v)$ Assume that $n$ is odd and that $H$ can either be alternating or symmetric. Note that, as the extension is split, $H$ defines two orbits of size $n$. Let $X$ be one of them. Then any $g\in G$ can be uniquely written as $ha$, for some $a\in A$ and for some $h\in H$. As $X^{ha}=X^{a}$, we need to find $a$ such that $X\cap X^{a}$ is empty. But, since every $a\in A$ swaps the two elements inside an even number of blocks, this is not possible. Hence there does not exist a $g\in G$ that such that $X\cap X^{g}$ is empty. In particular, $\mathbf{m}(G)\leq n$ and $G$ does not appear in the list. This concludes the proof, as all the transitive groups of $C_{2}\mathop{\mathrm{wr}}\mathrm{Sym}(n)$ that induces either $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$ on $\Sigma$. ∎
Suppose that $G$ is imprimitive with two blocks of cardinality $n$. Let us denote these blocks by $B$ and $C$ and let $X,Y$ be two subsets of $B$ of equal size $k<n/2$. We will show that there exists a permutation in $G$ mapping $X$ to $Y$. Let $g$ be an arbitrary permutation that swaps the two blocks $B$ and $C$. Consider the subset
$A=X\cup\left(C\setminus Y^{g^{-1}}\right).$
Since $|A|=n<\mathbf{m}(G)$, there exists a permutation $h$ such that $A\cap A^{h}$ is empty. Observe that $X^{h}=Y^{g^{-1}}$ and hence $X^{hg}=Y$. By the arbitrariness of $X$ and $Y$, we conclude that the permutation group that $G^{B}_{B}$, induced by the action of $G_{B}$ on $B$, is $k$-homogeneous, for every $k<n/2$. Using again [13, Section 5] and [18, Theorem 9.4B], we have that the remaining possibilities for $G_{B}^{B}$ are as follows:
  1. $(e)$
    $\mathrm{Alt}(n)\leq G_{B}^{B}\leq\mathrm{Sym}(n)$ in their natural action on $n$ points;
  2. $(f)$
    $\mathrm{ASL}_{d}(q)\leq G_{B}^{B}\leq\mathrm{A}\Gamma\mathrm{L}_{d}(q)$ in their natural action on $q^{d}$ points, where
    $(d,q)\in\{(1,4),(1,5),(1,8),(2,2),(3,2)\};$
  3. $(g)$
    $\mathrm{PSL}_{d}(q)\leq H\leq\mathrm{P}\Gamma\mathrm{L}_{2}(q)$ in their natural action on $\mathrm{PG}_{1}(q)$, where $q\in\{4,5,7,8\}$;
  4. $(h)$
    $G_{B}^{B}$ is isomorphic to $M_{12}$ in its natural actions on $12$ points.
Once again, we need to consider these cases in turn. We postpone cases (f)(f)$(g)$ and  $(h)$ to the section of the proof where we collect our algorithmical considerations. In contrast, we cannot handle case $(e)$ computationally. The structure of the groups satisfying all the hypothesis we imposed can be quite wild, but can have a good understanding of some relevant subgroups for our purposes.
Assume that $n\geq 5$, and let $G$ be an imprimitive permutation group with two blocks, $B$ and $C$, of size $n$ such that $\mathrm{Alt}(n)\leq G_{B}^{B}\leq\mathrm{Sym}(n)$. We compute the socle of $G$. Preliminary, we note that $G_{(\Sigma)}=G_{B}=G_{C}$ and that $G$ is an extension of $C_{2}$ by $G_{(\Sigma)}$. Let $T$ be a minimal normal subgroup of $G_{(\Sigma)}$. Consider the projections $\pi_{B}:G_{(\Sigma)}\to G_{B}^{B}$ and $\pi_{C}:G_{(\Sigma)}\to G_{C}^{C}$. Since $T$ is minimal normal, so are $T^{\pi_{B}}$ and $T^{\pi_{C}}$. Observe that, as $G$ is faithful, the two images cannot be concurrently trivial. In particular, $\mathrm{Alt}(n)$ is a subgroup of $T$, and hence the socle of $G$ is of the form $\mathrm{Alt}(n)^{\ell}$, for some positive integer $\ell$. As $G\leq\mathrm{Sym}(n)\mathop{\mathrm{wr}}C_{2}$, we see that $\ell\leq 2$, and hence either $\ell=1$ or $\ell=2$. If $\ell=2$, it follows that
$\mathrm{Alt}(n)\mathop{\mathrm{wr}}C_{2}\leq G\leq\mathrm{Sym}(n)\mathop{\mathrm{wr}}C_{2}.$
On the other hand, suppose that $\ell=1$. Since $G$ acts transitively on $\Sigma$, there exists a permutation $s\in G$ such that $B^{s}=C$ and $s^{2}\in G_{(\Sigma)}$. In particular, depending on whether $G_{B}^{B}$ is alternating or symmetric, $G$ contains a subgroup isomorphic to either $\mathrm{Alt}(n)\langle s\rangle$ or $\mathrm{Sym}(n)\langle s\rangle$. Observe that $G_{B}^{B}$ and $G_{C}^{C}$ are permutationally isomorphic: the requested bijections are $s$ for the domains and the conjugation by $s$ for the groups. By [18, Exercise 1.6.1], there exists $f\in\mathrm{Sym}(n)$ such that every element of $G$ can be written as $(x,x^{f})$, where $x$ is a permutation of $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$ in natural action. The possibilities for $G$ depend on whether the action induced on the two blocks are equivalent or not. By assuming that $n\geq 7$, and by applying, for instance, [18, Lemma 1.6B], we find that the alternating group has two (permutationally isomorphic but) not equivalent actions on $n$ points (and one is mapped to the other by applying an outer automorphism of $\mathrm{Alt}(n)$), while the symmetric group has only one. Therefore, there are three representatives for the permutational isomorphism classes for $G$.
  • If $G_{B}^{B}=\mathrm{Alt}(n)$ and $f$ is even, then $G$ is permutationally isomorphic to $\mathrm{Alt}(n)\times C_{2}$;
  • If $G_{B}^{B}=\mathrm{Alt}(n)$ and $f$ is odd, then $G$ is permutationally isomorphic to $\mathrm{Alt}(n)\rtimes\langle s\rangle$, where $s^{2}=1$ and $s$ applies a transposition while swapping the blocks;
  • If $G_{B}^{B}=\mathrm{Sym}(n)$, then $G$ is permutationally isomorphic to $\mathrm{Sym}(n)\times C_{2}$.
(We point out that, by repeating the same reasoning, we can describe the transitive permutation groups arising in case $(h)$. Indeed, as $N_{\mathrm{Sym}(12)}(M_{12})/M_{12}^{2}\cong\mathrm{Out}(M_{12})$ has order $2$, we find that $G$ is isomorphic to either $M_{12}\mathop{\mathrm{wr}}C_{2}$, to $M_{12}\times C_{2}$, or to $M_{12}\langle s\rangle$, where the $s$ swaps the blocks and the action that the subgroup $M_{12}$ induces on the two blocks are not equivalent. In our algorithm, we actually built these groups as the only three transitive subgroups of $M_{12}\mathop{\mathrm{wr}}C_{2}$, rather than using this theoretical approach.)
We can now establish which groups from $(e)$ meet the upper bound from D.
Lemma 28.
Let $n\geq 8$ be a positive integer, and let $G$ be an imprimitive group with block system $\Sigma$ that consists of two blocks of size $n$. Suppose that the action of $G$ on a block is permutationally isomorphic to $\mathrm{Alt}(n)$ or $\mathrm{Sym}(n)$. Then $\mathbf{m}(G)=n+1$.
Proof.
First, we have verified that the statement is true for $n=8$ with a computer-aided calculation. Hence, during the proof, we assume that $n\geq 9$. We split the discussion in two cases, depending on the socle of $G$ being isomorphic to $\mathrm{Alt}(n)^{2}$ or to $\mathrm{Alt}(n)$.
Suppose that $\mathrm{Alt}(n)\mathop{\mathrm{wr}}C_{2}\leq G$, and let $A$ be a $k$-subset of points, for some $k\leq n$. We write $k_{B}=|A\cap B|$ and $k_{C}=|A\cap C|$. If $\max\{k_{B},k_{C}\}\leq\mathbf{m}(\mathrm{Alt}(n))$, then there exists a permutation $g\in\mathrm{Alt}(n)^{2}$ such that $A\cap A^{g}$ is empty. Otherwise, as $\mathbf{m}(\mathrm{Alt}(n))>n/2$, we can assume (by eventually renaming the blocks) that
$k_{C}=k-k_{B}<n/2.$
Let $s\in C_{2}$ be the element of $G$ swapping the two blocks. Note that
$|(A\cap B)^{s}\cap(A\cap C)|\leq k_{C}<n/2\quad\hbox{and}\quad|(A\cap C)^{s}\cap(A\cap B)|\leq k_{C}<n/2.$
Thus, by $(n/2)$-transitivity of $\mathrm{Alt}(n)$, we can choose $(h_{B},h_{C})\in\mathrm{Alt}(n)^{2}$ such that
$(A\cap B)^{sh_{C}}\cap(A\cap C)=\emptyset\quad\hbox{and}\quad(A\cap C)^{sh_{B}}\cap(A\cap B)=\emptyset.$
Hence, $g=s(h_{B},h_{C})\in G$ has the property that $A$ and $A^{g}$ are disjoint. This exhausts the first case.
Suppose that $\mathrm{Alt}(n)\langle s\rangle\leq G$ but $\mathrm{Alt}(n)\mathop{\mathrm{wr}}C_{2}$ is not a subgroup of $G$. If $s$ is central in $\mathrm{Alt}(n)\langle s\rangle$, $G$ has blocks of size two, and hence the conclusion follows from Lemma 27. Assume now that $s$ is not central in $\mathrm{Alt}(n)\langle s\rangle$. Let $A$ be a $k$-subset of points, for some $4\leq k\leq n$. We can suppose that $A$ and $A^{s}$ are not disjoint. We label the points of $B$ by $\{1_{B},2_{B},\dots,n_{B}\}$, and those of $C$ by $\{1_{C},2_{C},\dots,n_{C}\}$. By the discussion preceding this lemma, we can also assume that $1_{B}^{s}=2_{C}$, $2_{B}^{s}=1_{C}$, $1_{C}^{s}=2_{B}$, $2_{C}^{s}=1_{B}$, while, for every integer $3\leq x\leq n$, $x_{B}^{s}=x_{C}$ and $x_{C}^{s}=x_{B}$. We define the integers
$z=|A\cap\left\{1_{B},2_{B},1_{C},2_{C}\right\}|,\quad z_{B}=|A\cap\left\{1_{B},2_{B}\right\}|,\quad z_{C}=|A\cap\left\{1_{C},2_{C}\right\}|.$
Observe that there exist two integers $a,b\leq n$ such that
$4-z=|A\cap\left\{a_{B},b_{B},a_{C},b_{C}\right\}|,\quad 2-z_{B}=|A\cap\left\{a_{B},b_{B}\right\}|,\quad 2-z_{C}=|A\cap\left\{a_{C},b_{C}\right\}|.$
It is immediate to check that there exists a double transposition $t$ such that $t\in\mathrm{Alt}(4)$ and
$\left\{1_{B},2_{B},1_{C},2_{C},a_{B},b_{B},a_{C},b_{C}\right\}\cap\left\{1_{B},2_{B},1_{C},2_{C},a_{B},b_{B},a_{C},b_{C}\right\}^{st}=\emptyset.$
Finally, the pointwise stabilizer of the set $\Omega-\left\{1_{B},2_{B},1_{C},2_{C},a_{B},b_{B},a_{C},b_{C}\right\}$ contains $\mathrm{Alt}(n-4)$. The choice of a suitable $h$ so that $A$ and $A^{sth}$ are disjoint is the same as in the corresponding case in Lemma 27. (Note that our proof of the equality $\mathbf{m}(\mathrm{Alt}(n)\times C_{2})=\mathbf{m}(\mathrm{Sym}(n)\times C_{2})=n+1$ does not require $n\geq 8$, but rather the weaker $n\geq 5$: this is what forces us to take $n\geq 9$ in this proof.)
Since no more transitive groups with the required properties exist, the proof is complete. ∎
We now focus on establishing which of the listed groups meet the putative upper bound. To understand which possibility in cases (b)(b)(c)(c)(f)(f)$(g)$ and $(h)$ achieve the upper bound, we run a computer-aided calculation, whose code is available at [5]. Initially, we filtered the imprimitive permutation groups of degree up to $18$ with the property that they define a block system $\Sigma$ with $G_{B}^{B}$ and $G^{\Sigma}$ as prescribed. Then, we applied a greedy approach to establish if they achieve or not the upper bound by checking whether there exists a $n$-subset of the domain that is not self-separable. (We point out that, for this algorithm, we have not used Lemma 23 – we are unaware if filtering these groups using this criteria would speed up the computation.)
This line of attack seems to be too time consuming for case $(d)$. Running the code on the single group $C_{2}\mathop{\mathrm{wr}}M_{11}$ shows that $\mathbf{m}(C_{2}\mathop{\mathrm{wr}}M_{11})\leq 11$, and hence the putative upper bound of $12$ is not reached. In particular, by Lemma 12, this excludes that any subgroup of $C_{2}\mathop{\mathrm{wr}}M_{11}$ meets the upper bound. On the other hand, $\mathbf{m}(C_{2}\mathop{\mathrm{wr}}M_{12})=13$. We should not establish which subgroup of the wreath product is transitive on $24$ points: this task is computationally too demeaning, and this is the reason why earlier we described the transitive subgroups of $C_{2}\mathop{\mathrm{wr}}M_{12}$ that induce $M_{12}$ as their permutation group on the block system. In particular, we feed these transitive groups to the algorithm that checks whether subsets of $12$-points are self-separable. We obtain that all of them are for all the groups under consideration, with the exception of the perfect central extension of $C_{2}$ by $M_{12}$. We have thus considered all the group arising in $(d)$.
Proof of Lemma 22.
The statement for permutation groups whose degree is at most $24$ follows from the preceding discussion and the computation of [5]. The remaining possibilities has been addressed in Lemma 27 and Lemma 28. No other case need to be considered, and thus our proof is complete. ∎
8. Primitive groups
This section is devoted to the proofs of our results concerning primitive permutation groups. The subsections are organized according to the type of primitive group under consideration, as classified by the O’Nan–Scott Theorem. As previously mentioned, we follow the partition of primitive groups into eight classes proposed in [38].
8.1. Almost simple type
The proof of J is divided into several lemmas depending on the socle of the group considered.
We start by considering the action of the alternating group $\mathrm{Alt}(m)$ or the symmetric group $\mathrm{Sym}(m)$ on the set $m]\choose k$} of $k$-subsets of the set $m]=\{1,2,\ldots,m\}$. Observe that the degree of this permutation representation is, $m\choose k$}. (Lemma 29 takes care of the case $(a)$ in Table 1.)
Lemma 29.
Let $2\leq k<m$ be two integers. Then
[…]
Consequently, if $\mathcal{C}_{k}$ is the family of symmetric (or alternating) groups acting on $k$-subsets, then
$\lim_{n}\frac{\mathbf{F}_{\mathcal{C}_{k}}\left(n\right)}{\sqrt{n}}\leq\frac{\sqrt{k!}}{(k/2)!}\quad\hbox{ for }k\hbox{ even,}$
and
$\lim_{n}\frac{\mathbf{F}_{\mathcal{C}_{k}}\left(n\right)}{n^{\frac{k+1}{2k}}}\leq\frac{k!^{\frac{k+1}{2k}}}{\lceil k/2\rceil!}\quad\hbox{ for }k\hbox{ odd}.$
Proof.
For the sake of simplicity, let $r=\lfloor k/2\rfloor$. Consider the subset
$A=\left\{X\in{[m]\choose k}\,\middle|\,\{1,2,\ldots,r\}\subseteq X\right\}.$
For every $g\in\mathrm{Sym}(m)$, we compute
$A\cap A^{g}=\left\{X\in{[m]\choose k}\,\middle|\,\{1,2,\ldots,r\}\cup\{1^{g},2^{g},\ldots,r^{g}\}\subseteq X\right\}.$
This intersection is never empty, and, if $g$ does not stabilise $\{1,2,\ldots,r\}$, $A^{g}$ and $A$ are distinct. Hence, $A$ is not self-separable for $\mathrm{Sym}(m)$. This considerations prove the upper bound of $\mathbf{m}(\mathrm{Sym}(m))$.
Note that
$|A|={m-r\choose k-r}=\frac{1}{(k-r)!}m^{k-r}+o\left(m^{k-r}\right),\quad\hbox{and}\quad n={m\choose k}=\frac{1}{k!}m^{k}+o\left(m^{k}\right)\,.$
A direct computation is now sufficient to determine the claimed asymptotic behaviours of $\mathbf{f}_{\mathcal{C}}(n)$ and $\mathbf{F}_{\mathcal{C}}(n)$. ∎
We now turn our attention to the standard actions of the classical groups. We are going to follow the notation established in [11, Chapters 2 and 4].
We start by considering the possible subspace actions of a classical group whose socle is isomorphic to $\mathrm{PSL}_{d}(q)$. There are three actions to consider: that on $k$-subspaces, that on pairs of complementing subspaces, and that on pairs of subspaces, one containing the other. From here on, the symbol $V^{(k)}$ denotes the set of $k$-dimensional subspaces of the natural module of a classical group. (Lemma 30 takes care of the cases $(b)$, $(c)$ and $(d)$ in Table 1.)
Lemma 30.
Let $G$ be a classical group with linear socle and associated natural module $\mathbb{F}_{q}^{d}$ endowed with a subspace action.
  1. $(i)$
    If the domain of the action is the set of $k$-subsets, with $k\leq d/2$, then
    $\mathbf{m}(G)\leq\left[d-\lfloor k/2\rfloor\atop k-\lfloor k/2\rfloor\right]_{q}.$
    Consequently, if $\mathcal{C}_{k}$ is the family of such permutation groups, then
    $\lim_{n}\frac{\mathbf{f}_{\mathcal{C}_{k}}\left(n\right)}{\sqrt{n}}=\lim_{n}\frac{\mathbf{F}_{\mathcal{C}_{k}}\left(n\right)}{\sqrt{n}}=1\quad\hbox{ for }k\hbox{ even,}$
    and
    $\lim_{n}\frac{\mathbf{F}_{\mathcal{C}_{k}}\left(n\right)}{n^{\frac{k+1}{2k}}}\leq 1\quad\hbox{ for }k\hbox{ odd}.$
  2. $(ii)$
    If the domain of the action is the set of pairs of subspaces complementing each other, one of which of dimension $k\leq d/2$, then
    $\mathbf{m}(G)\leq q^{k(d-k)}\left[d-\lfloor k/2\rfloor\atop k-\lfloor k/2\rfloor\right]_{q}.$
    Consequently, if $\mathcal{C}_{k}$ is the family of such permutation groups, then
    $\lim_{n}\frac{\mathbf{f}_{\mathcal{C}_{k}}\left(n\right)}{\sqrt{n}}=\lim_{n}\frac{\mathbf{F}_{\mathcal{C}_{k}}\left(n\right)}{\sqrt{n}}=1\quad\hbox{ for }k\hbox{ even,}$
    and
    $\lim_{n}\frac{\mathbf{F}_{\mathcal{C}_{k}}\left(n\right)}{n^{\frac{k+1}{2k}}}\leq 1\quad\hbox{ for }k\hbox{ odd}.$
  3. $(iii)$
    If the domain of the action is the set of pairs of subspaces, one of dimension $k<d/2$ and the other of dimension dkd-k, the smaller contained in the larger, then
    $\mathbf{m}(G)\leq\left[\lceil(d-k)/2\rceil\atop\lceil(d-3k)/2\rceil\right]_{q}\left[\lceil(2d-3k)/2\rceil\atop\lceil k/2\rceil\right]_{q}.$
    Consequently, if $\mathcal{C}_{k}$ is the family of such permutation groups, then
    $\lim_{n}\frac{\mathbf{F}_{\mathcal{C}_{k}}\left(n\right)}{n^{\frac{k\left\lceil\frac{d-3k}{2}\right\rceil+(d-k)\left\lceil\frac{k}{2}\right\rceil}{k(2d-3k)}}}\leq 1.$
Proof.
We start by considering the action of $G$ on $k$-subspaces. Let
$U=\langle e_{1},e_{2},\ldots,e_{\lfloor k/2\rfloor}\rangle,$
and consider the subset
$A=\left\{X\in V^{(k)}\,\middle|\,U\leq X\right\}.$
For every $g\in G$, we can choose a subspace $X_{g}\in V^{(k)}$ such that
$U+U^{g^{-1}}\leq X_{g}.$
(Note that such a subspace exists because $\dim(U+U^{g^{-1}})\leq k$.) By construction,
$U^{g}+U\leq X_{g}^{g}\in A\cap A^{g}.$
Hence $A\cap A^{g}$ is nonempty, and $A$ is self-separable. Observe that
$|A|=\left[d-\lfloor k/2\rfloor\atop k-\lfloor k/2\rfloor\right]_{q}=q^{\left(k-\left\lceil\frac{k}{2}\right\rceil\right)(d-k)}+O\left(q^{\left(k-\left\lceil\frac{k}{2}\right\rceil\right)(d-k)-1}\right),$
and
$|V^{(k)}|=\left[d\atop k\right]_{q}=q^{k(d-k)}+O\left(q^{k(d-k)-1}\right).$
Therefore, we can compute that
$|A|\sim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}.$
This concludes the proof of $(i)$.
Let us now consider the action on pairs of complementary subspaces. Suppose that the smallest one has dimension $k$. Define
$U=\langle e_{1},e_{2},\ldots,e_{\lfloor k/2\rfloor}\rangle,$
and consider
$A=\left\{(X,Y)\in V^{(k)}\times V^{(d-k)}\,\middle|\,V=X\oplus Y\hbox{ and }U\leq Y\right\}.$
For every $g\in G$, we choose $(X_{g},Y_{g})\in V^{(k)}\times V^{(d-k)}$ so that
$U+U^{g^{-1}}\leq X_{g}\quad\hbox{and}\quad V=X_{g}\oplus Y_{g}.$
A similar argument as in the first case shows that $A$ is not self-separable, because $(X_{g},Y_{g})^{g}\in A\cap A^{g}$. Moreover,
$|A|=q^{k(d-k)}\left[d-\lfloor k/2\rfloor\atop k-\lfloor k/2\rfloor\right]_{q}=q^{k(d-k)\left(k-\left\lceil\frac{k}{2}\right\rceil\right)(d-k)}+O\left(q^{k(d-k)\left(k-\left\lceil\frac{k}{2}\right\rceil\right)(d-k)-1}\right),$
and
$|\Omega|=q^{k(d-k)}\left[d\atop k\right]_{q}=q^{2k(d-k)}+O\left(q^{2k(d-k)-1}\right).$
Therefore,
$|A|\sim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil},$
which is enough to conclude $(ii)$.
Finally, we need to consider the domain of pairs of subspaces, say $X$ and $Y$, so that $X\in V^{(k)}$, $Y\in V^{(d-k)}$ and $X\leq Y$. In this case, we have to set
$U=\langle e_{1},\ldots,e_{\lfloor k/2\rfloor}\rangle$
and
$W=\langle e_{1},\ldots,e_{\lfloor(d-k)/2\rfloor}\rangle.$
Considering the set
$A=\{(X,Y)\in V^{(k)}\times V^{(d-k)}\mid X\leq Y,\,U\leq X,\,W\leq Y\},$
and using similar strategies as before, we conclude that $A$ is not self-separable. In particular, we have
$|A|$ $=\left[d-\lfloor(d-k)/2\rfloor\atop d-k-\lfloor(d-k)/2\rfloor\right]_{q}\left[d-k-\lfloor k/2\rfloor\atop k-\lfloor k/2\rfloor\right]_{q}$
$=\left[\lceil(d-k)/2\rceil\atop\lceil(d-3k)/2\rceil\right]_{q}\left[\lceil(2d-3k)/2\rceil\atop\lceil k/2\rceil\right]_{q}$
$=q^{k\left\lceil\frac{d-3k}{2}\right\rceil+(d-k)\left\lceil\frac{k}{2}\right\rceil}+O\left(q^{k\left\lceil\frac{d-3k}{2}\right\rceil+(d-k)\left\lceil\frac{k}{2}\right\rceil-1}\right)$
and
$|\Omega|=\left[d-k\atop k\right]_{q}\left[d\atop d-k\right]_{q}=q^{k(2d-3k)}+O\left(q^{k(2d-3k)-1}\right).$
Therefore,
$|A|\sim n^{\frac{k\left\lceil\frac{d-3k}{2}\right\rceil+(d-k)\left\lceil\frac{k}{2}\right\rceil}{k(2d-3k)}}.$
This completes the proof of the missing case $(iii)$. ∎
Let now $G$ be a classical group whose socle is not isomorphic to $\mathrm{PSL}_{d}(q)$. We start by considering the action of $G$ on the totally isotropic or totally singular $k$-subspaces of its natural module. From here on, with the symbol $V_{\bullet}^{(k)}$ we refer to the set of $k$-dimensional subspaces of the formed space $(V,\mathbf{b})$ where $\bullet$ signals a given property, which is clear from context. For instance, in the next proof $V_{\bullet}^{(k)}$ denotes in a unified way the totally isotropic and totally singular $k$-subspaces. (Lemma 31 takes care of the cases $(e)$, $(g)$, $(i)$, $(j)$ and $(k)$ in Table 1. Moreover, not to overcomplicate our formulae, we adopt the notation that if, in a product $\prod_{i=a}^{b}x_{i}$, $a$ is greater than $b$, then the product is equal to $1$.)
Lemma 31.
Let $G$ be a classical group whose socle is not linear, and consider its subspace action on totally isotropic or totally singular $k$-subspaces of its natural module $V$. Then there exists $A$ which is a non self-separable subset of the domain whose exact cardinality, as well as the cardinality of the domain, are collected in Table 3.
Observe that in the first row of Table 3 the symbol $\delta$ appears. We have that $\delta=|(2,d)-2|$, that is, $\delta=0$ if $d$ is even, and $\delta=1$ if $d$ is odd.
Moreover, the expression of $|\Omega|$ and $|A|$ appearing in Table 3 appear to be to complicated to obtain an explicit asymptotic behaviour. We shall be satisfied by observing that the upper bound that we have computed can be (trivially) formulated as
$|\Omega|^{\log_{|\Omega|}|A|},$
and that many experimentation suggests that, for a fixed $k$, the exponent should be arbitrarily close to $1/2$ if we allow $d$ to grow to infinity.
Proof of Lemma 31.
For the sake of uniformity, let us introduce the parameter
$m=\dim(V)-2\dim(M),$
where $M$ is any totally isotropic or or totally singular subspace of maximal dimension. We fix a totally isotropic or a totally singular subspace $U$ with
$\dim(U)=\frac{\dim(V)-m}{2}-k.$
Consider the set
$A=\left\{X\in V^{(k)}_{\bullet}\mid X\perp U\right\},$
and let $M$ be a totally isotropic or totally singular subspace of maximal dimension containing $U$. For every $g\in G$, we define
$X_{g}=(U+U^{g^{-1}})^{\perp}\cap M.$
Note that
$\dim(X_{g})=\dim(M)-\dim\left(\frac{U^{g^{-1}}}{U^{g^{-1}}\cap M}\right)\geq\frac{\dim(V)-m}{2}-\dim(U)=k.$
We can now choose $Y^{g}$, a $k$-subspace of $X_{g}$. Observe that
$Y_{g}\leq X_{g}\leq M=M^{\perp}\leq X_{g}^{\perp}\leq Y_{g}^{\perp}$
and that, by construction,
$Y_{g}^{g}\perp U^{g}+U,$
and thus $Y_{g}^{g}\in A\cap A^{g}$. Therefore, $A$ is not self-separable.
We now need to focus to compute $|A|$ and the degree of the actions. Note that the latter is the number of totally isotropic or total singular $k$-subspaces in a given formed space: this value can be found in [11, Table 4.1.2]. On the other hand, to count the elements of $A$ we just need to count the number of totally isotropic or totally singular subspaces of $U^{\perp}$. Observe that, as $\mathbf{b}\downarrow_{U}$ is trivial, $U^{\perp}$ can be written as the orthogonal direct sum $U^{\perp}=U\oplus U^{\perp}/U$, where $U$ is totally isotropic or totally singular, and $(U^{\perp}/U,\mathbf{b}\downarrow_{U^{\perp}/U})$ is a formed space of the same type as $(V,\mathbf{b})$. Furthermore,
$\dim(U^{\perp})=\dim(V)-\dim(M)+k=\frac{\dim(V)+m}{2}+k,$
and
$\dim(U^{\perp}/U)=\dim(V)-2\dim(M)+2k=2k+m.$
To every $X\in A$, we can uniquely assign a direct sum decomposition
$X=\left(X\cap U\right)\oplus\left(X\cap U^{\perp}/U\right).$
Hence, $H$ can be identified with a pair $(X_{0},X_{1})$, where $X_{1}$ is a totally isotropic or totally singular $h$-subspace in $U^{\perp}/U$, and $X_{0}$ with a (kh)(k-h)-subspace of $U$. Therefore, if $N_{h}(U^{\perp}/U)$ denotes the number of totally isotropic or totally singular $h$-subspaces of $U^{\perp}/U$, we proved that
$|A|=\sum_{h=\max\{0,2k-\dim(M)\}}^{k}\left[\dim(M)-k\atop k-h\right]_{q}N_{h}(U^{\perp}/U).$
(Note that the Gaussian coefficient loses meaning if $\dim(M)-k<k-h$. This explains why $h$ varies in the prescribed range.) Finally, by using the formulae of [11, Table 4.1.2], we can compute $|A|$. We have explained a way to compute all the data appearing in Table 3, which completes the proof of Lemma 31. ∎
$\mathrm{soc}(G)$ $|\Omega|$ $|A|$
$(i)$ $\mathrm{PSU}_{d}(q)$ $\frac{\prod\limits_{i=d-2k+1}^{d}(q^{i}-(-1)^{i})}{\prod\limits_{i=1}^{k}(q^{2i}-1)}$ $\sum\limits_{h=\max\{0,2k-\lfloor\frac{d}{2}\rfloor\}}^{k}\left[\lfloor\frac{d}{2}\rfloor-k\atop k-h\right]_{q}\frac{\prod\limits_{i=2k-2h+1+\delta}^{2k+\delta}(q^{i}-(-1)^{i})}{\prod\limits_{i=1}^{h}(q^{2i}-1)}$
$(ii)$ $\mathrm{PSp}_{2d}(q)$ $\frac{\prod\limits_{i=d-k+1}^{d}(q^{2i}-1)}{\prod\limits_{i=1}^{k}(q^{i}-1)}$ $\sum\limits_{h=\max\{0,2k-d\}}^{k}\left[d-k\atop k-h\right]_{q}\frac{\prod\limits_{i=k-h+1}^{k}(q^{2i}-1)}{\prod\limits_{i=1}^{h}(q^{i}-1)}$
$(iii)$ $\mathrm{P}\Omega_{2d}^{+}(q)$ $\frac{(q^{d}-1)(q^{d-k}+1)\prod\limits_{i=d-k+1}^{d-1}(q^{2i}-1)}{2^{\delta(d,k)}\prod\limits_{i=1}^{k}(q^{i}-1)}$ $\sum\limits_{h=\max\{0,2k-d\}}^{k}\left[d-k\atop k-h\right]_{q}\frac{(q^{k}-1)(q^{k-h}+1)\prod\limits_{i=k-h+1}^{k-1}(q^{2i}-1)}{2^{\delta(k,h)}\prod\limits_{i=1}^{h}(q^{i}-1)}$
$(iv)$ $\mathrm{P}\Omega_{2d}^{-}(q)$ $\frac{(q^{d}+1)(q^{d-k}-1)\prod\limits_{i=d-k+1}^{d-1}(q^{2i}-1)}{\prod\limits_{i=1}^{k}(q^{i}-1)}$ $\sum\limits_{h=\max\{0,2k-d-1\}}^{k}\left[d-1-k\atop k-h\right]_{q}\frac{(q^{k+1}+1)(q^{k-h+1}-1)\prod\limits_{i=k-h+2}^{k}(q^{2i}-1)}{\prod\limits_{i=1}^{h}(q^{i}-1)}$
$(v)$ $\Omega_{2d+1}(q)$ $\frac{\prod\limits_{i=d-k+1}^{d}(q^{2i}-1)}{\prod\limits_{i=1}^{k}(q^{i}-1)}$ $\sum\limits_{h=\max\{0,2k-d\rfloor\}}^{k}\left[d-k\atop k-h\right]_{q}\frac{\prod\limits_{i=k-h+1}^{k}(q^{2i}-1)}{\prod\limits_{i=1}^{h}(q^{i}-1)}$
Table 3. Computation of degree, $|A|$, and asymptotics for $|A|$ for classical groups acting on totally singular or totally isotropic $k$-subspaces.
Once again, let $G$ be a classical group whose socle is not isomorphic to $\mathrm{PSL}_{d}(q)$. We study the action of $G$ on the nondegenerate $k$-subspaces of its natural module. (Lemma 32 takes care of the cases $(f)$, $(h)$, $(m)$, $(n)$, $(o)$, $(p)$, $(q)$, $(r)$ and $(s)$ in Table 1.)
Lemma 32.
Let $G$ be a classical group whose socle is not linear, and consider its subspace action on nondegenerate $k$-subspaces of its natural module $V$. Then there exists $A$ which is a non self-separable subset of the domain whose exact cardinality and asymptotic size in relation with the degree are collected in Table 4.
In Table 4, the letters $B$ and $C$ replace the functions
$B(k,d)=-2+3d-2k\quad\hbox{and}\quad C(k,d)=-2+4d-4k.$
In particular, it is interesting for note that, for $k$ fixed,
$\lim_{d}\frac{B(k,d)}{C(k,d)}=\frac{3}{4}\quad\hbox{and}\quad\lim_{d}\frac{1}{C(k,d)}=0.$
Proof.
As in the proof of Lemma 31, we start by fixing a subspace $U$. Unlike before, we need to make from the start different choices of $U$ depending on the socle of $G$ and, if the bilinear form is symmetric, depending if the subspaces we are acting on are hyperbolic, parabolic or elliptic. These choices are as follows.
  1. $(i)$
    If $\mathrm{soc}(G)$ is unitary, and we are acting on nondegenerate $k$-subspaces, $U$ is any nondegenerate $\lfloor k/2\rfloor$-subspace.
  2. $(ii)$
    If $\mathrm{soc}(G)$ is symplectic, and we are acting on nondegenerate $2k$-subspaces, $U$ is any nondegenerate $2\lfloor k/2\rfloor$-subspace.
  3. $(iii)$
    If $\mathrm{soc}(G)$ is orthogonal, and we are acting on nondegenerate hyperbolic $2k$-subspaces, $U$ is any nondegenerate hyperbolic $2\lfloor k/2\rfloor$-subspace.
  4. $(iv)$
    If $\mathrm{soc}(G)$ is orthogonal, and we are acting on nondegenerate parabolic $(2k+1)$-subspaces, $U$ is any nondegenerate hyperbolic $2\lfloor k/2\rfloor$-subspace.
  5. $(v)$
    If $\mathrm{soc}(G)$ is orthogonal, and we are acting on nondegenerate elliptic $2k$-subspaces, $U$ is any nondegenerate hyperbolic $2\lfloor(k-1)/2\rfloor$-subspace.
Consider
$A=\left\{X\in V^{(h)}_{\bullet}\mid U\leq X\right\},$
where $h\in\{k,2k,2k+1\}$ depending on the case we are considering. For every $g\in G$, define
$X_{g}=U+U^{g^{-1}}.$
Note that the restricted form $\mathbf{b}\downarrow_{X_{g}}$ is sequilinear, alternating or symmetric (hyperbolic or parabolic depending on the dimension of $X$) according to the socle being unitary, symplectic or orthogonal, respectively. Furthermore,
$\dim(X_{g})\leq 2\dim(U)\leq 2^{\epsilon}k-\delta,$
where $\epsilon=0$ if the socle is unitary, and $\epsilon=1$ otherwise, and $\delta=1$ if the socle is orthogonal and the formed subspace is parabolic, $\delta=2$ if the socle is orthogonal and the formed subspace is elliptic, and $\delta=0$ otherwise. It is now immediate to build a subspace $Y_{g}\in V^{(k)}_{\bullet}$ such that $X_{g}\leq Y_{g}$. Moreover, by construction, $Y_{g}\in A\cap A^{g}$. Therefore, $A$ is not self-separable.
We need to deal with computing the size of $A$. Observe that, in any case, $(U,\mathbf{b}\downarrow_{U})$ is a nondegenerate formed space whose bilinear form is either sesquilinear, alternating or hyperbolic bilinear, depending on the socle of $G$. Therefore, by quotienting by $U$, AA can be thought of as the set of nondegenerate $(h-\dim(U))$-subspaces of $V/U$. Therefore, the formulae for computing $|A|$ and the degree are readily available in [11, Table 4.1.2]. This information completes the proof of Lemma 32.
$\mathrm{soc}(G)$ Domain $|\Omega|$ $|A|$ Asymptotics
$(i)$ $\mathrm{PSU}_{d}(q)$ $k$-nondegenerate $\frac{q^{k(d-k)}\prod\limits_{i=d-k+1}^{d}(q^{i}-(-1)^{i})}{\prod\limits_{i=1}^{k}(q^{i}-(-1)^{i})}$ $\frac{q^{\lceil\frac{k}{2}\rceil(d-k)}\prod\limits_{i=d-k+1}^{d-\lfloor\frac{k}{2}\rfloor}(q^{i}-(-1)^{i})}{\prod\limits_{i=1}^{\lceil\frac{k}{2}\rceil}(q^{i}-(-1)^{i})}$ $\lesssim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(ii)$ $\mathrm{PSp}_{2d}(q)$ $2k$-nondegenerate $\frac{q^{2k(d-k)}\prod\limits_{i=d-k+1}^{d}(q^{2i}-1)}{\prod\limits_{i=1}^{k}(q^{2i}-1)}$ $\frac{q^{2\lceil\frac{k}{2}\rceil(d-k)}\prod\limits_{i=d-k+1}^{d-\lfloor\frac{k}{2}\rfloor}(q^{2i}-1)}{\prod\limits_{i=1}^{\lceil\frac{k}{2}\rceil}(q^{2i}-1)}$ $\lesssim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(iii)$ $\mathrm{P}\Omega_{2d}^{+}(q)$ $2k$-hyperbolic $\frac{q^{2k(d-k)}(q^{d}-1)\prod\limits_{i=d-k}^{d-1}(q^{2i}-1)}{2(q^{k}-1)(q^{d-k}-1)\prod\limits_{i=1}^{k-1}(q^{2i}-1)}$ $\frac{q^{2\lceil\frac{k}{2}\rceil(d-k)}(q^{d-\lfloor\frac{k}{2}\rfloor}-1)\prod\limits_{i=d-k}^{d-\lfloor\frac{k}{2}\rfloor-1}(q^{2i}-1)}{2(q^{\lceil\frac{k}{2}\rceil}-1)(q^{d-k}-1)\prod\limits_{i=1}^{\lceil\frac{k}{2}\rceil-1}(q^{2i}-1)}$ $\lesssim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(iv)$ $\mathrm{P}\Omega_{2d}^{+}(q)$ $(2k+1)$-parabolic $\frac{q^{(2k+1)(2d-2k-1)-1}(q^{d}-1)\prod\limits_{i=d-k}^{d-1}(q^{2i}-1)}{2\prod\limits_{i=1}^{k}(q^{2i}-1)}$ $\frac{q^{(2\lceil\frac{k}{2}\rceil+1)(2d-2k-1)-1}(q^{d-\lfloor\frac{k}{2}\rfloor}-1)\prod\limits_{i=d-k}^{d-\lfloor\frac{k}{2}\rfloor-1}(q^{2i}-1)}{2\prod\limits_{i=1}^{\lceil\frac{k}{2}\rceil}(q^{2i}-1)}$ $\lesssim n^{\frac{\lceil\frac{k}{2}\rceil+\frac{B}{C}-\frac{1}{C}\lfloor\frac{k}{2}\rfloor}{k+\frac{B}{C}}}$
$(v)$ $\mathrm{P}\Omega_{2d}^{+}(q)$ $2k$-elliptic $\frac{q^{2k(d-k)}(q^{d}-1)\prod\limits_{i=d-k}^{d-1}(q^{2i}-1)}{2(q^{k}+1)(q^{d-k}+1)\prod\limits_{i=1}^{k-1}(q^{2i}-1)}$ $\frac{q^{2\lceil\frac{k+1}{2}\rceil(d-k)}(q^{d-\lfloor\frac{k-1}{2}\rfloor}-1)\prod\limits_{i=d-k}^{d-\lfloor\frac{k-1}{2}\rfloor-1}(q^{2i}-1)}{2(q^{\lceil\frac{k+1}{2}\rceil}+1)(q^{d-k}+1)\prod\limits_{i=1}^{\lceil\frac{k-1}{2}\rceil}(q^{2i}-1)}$ $\lesssim n^{\frac{1}{k}\left\lceil\frac{k+1}{2}\right\rceil}$
$(vi)$ $\mathrm{P}\Omega_{2d}^{-}(q)$ $(2k+1)$-parabolic $\frac{q^{(2k+1)(2d-2k-1)-1}(q^{d}+1)\prod\limits_{i=d-k}^{d-1}(q^{2i}-1)}{2\prod\limits_{i=1}^{k}(q^{2i}-1)}$ $\frac{q^{(2\lceil\frac{k}{2}\rceil+1)(2d-2k-1)-1}(q^{d-\lfloor\frac{k}{2}\rfloor}+1)\prod\limits_{i=d-k}^{d-\lfloor\frac{k}{2}\rfloor-1}(q^{2i}-1)}{2\prod\limits_{i=1}^{\lceil\frac{k}{2}\rceil}(q^{2i}-1)}$ $\lesssim n^{\frac{\lceil\frac{k}{2}\rceil+\frac{B}{C}-\frac{1}{C}\lfloor\frac{k}{2}\rfloor}{k+\frac{B}{C}}}$
$(vii)$ $\mathrm{P}\Omega_{2d}^{-}(q)$ $2k$-elliptic $\frac{q^{2k(d-k)}(q^{d}+1)\prod\limits_{i=d-k}^{d-1}(q^{2i}-1)}{2(q^{k}+1)(q^{d-k}-1)\prod\limits_{i=1}^{k-1}(q^{2i}-1)}$ $\frac{q^{2\lceil\frac{k+1}{2}\rceil(d-k)}(q^{d-\lfloor\frac{k-1}{2}\rfloor}+1)\prod\limits_{i=d-k}^{d\lfloor\frac{k-1}{2}\rfloor-1}(q^{2i}-1)}{2(q^{\lceil\frac{k+1}{2}\rceil}+1)(q^{d-k}-1)\prod\limits_{i=1}^{\lceil\frac{k-1}{2}\rceil}(q^{2i}-1)}$ $\lesssim n^{\frac{1}{k}\left\lceil\frac{k+1}{2}\right\rceil}$
$(viii)$ $\Omega_{2d+1}(q)$ $2k$-hyperbolic $\frac{q^{k(2d-2k+1)}\prod\limits_{i=d-k+1}^{d}(q^{2i}-1)}{2(q^{k}-1)\prod\limits_{i=1}^{k-1}(q^{2i}-1)}$ $\frac{q^{\lceil\frac{k}{2}\rceil(2d-2k+1)}\prod\limits_{i=d-k+1}^{d-\lfloor\frac{k}{2}\rfloor}(q^{2i}-1)}{2(q^{\lceil\frac{k}{2}\rceil}-1)\prod\limits_{i=1}^{\lceil\frac{k}{2}\rceil-1}(q^{2i}-1)}$ $\lesssim n^{\frac{1}{k}\left\lceil\frac{k}{2}\right\rceil}$
$(ix)$ $\Omega_{2d+1}(q)$ $2k$-elliptic $\frac{q^{k(2d+1-2k)}\prod\limits_{i=d-k+1}^{d}(q^{2i}-1)}{2(q^{k}+1)\prod\limits_{i=1}^{k-1}(q^{2i}-1)}$ $\frac{q^{\lceil\frac{k+1}{2}\rceil(2d-2k+1)}\prod\limits_{i=d-k+1}^{d-\lfloor\frac{k-1}{2}\rfloor}(q^{2i}-1)}{2(q^{\lceil\frac{k+1}{2}\rceil}+1)\prod\limits_{i=1}^{\lceil\frac{k-1}{2}\rceil}(q^{2i}-1)}$ $\lesssim n^{\frac{1}{k}\left\lceil\frac{k+1}{2}\right\rceil}$
Table 4. Computation of degree, $|A|$, and asymptotics for $|A|$ for classical groups acting on nondegenerate $h$-subspaces.
We end Section 8.1 on a sour note. An attentive reader might have noticed that we have not discussed case $(l)$ in Table 1, which is the last step to complete the proof of J. In this case, the socle of $G$ is isomorphic to $\Omega^{\epsilon}_{2d}(2^{f})$, and the domain consists of nonsingular $1$-subspaces of the natural module of $G$. As stated in Remark 11, our current method does not give us any control on $1$-dimensional subspaces, and we have not developed an ad hoc approach for this scenario. Hence, the bound that appears in J is actually the general bound from D.
8.2. Simple diagonal type
Let us first recall the definition of primitive groups of simple diagonal type. Let $T$ be a nonabelian simple group, and let $k\geq 3$ be an integer. Note that we can naturally define three actions on $T^{k}$ – the right-regular permutation representation of $T^{k}$, the component-wise action of $\mathrm{Aut}(T)$, and the action of $\mathrm{Sym}(k)$ permuting indices. All these actions preserve the diagonal of $T^{k}$,
$\mathrm{diag}(T^{k})=\{(t,\ldots,t)\mid t\in T\}\,$,
hence we can consider their action on the domain $\Omega=T^{k}/\mathrm{diag}(T^{k})$. For simplicity, we can identify $\Omega$ with $T^{k-1}$ via the set of representatives
$\{(t_{1},\ldots,t_{k-1},1)\mid t_{1},\ldots,t_{k-1}\in T\}\,.$
Observe that, via this identification, the first k1k-1 components of $T^{k}$ act on $\Omega$ regularly by right multiplication, while the last component acts semiregularly by left multiplication. In particular, every inner automorphism can be realized by choosing two suitable elements in $T^{k-1}$ and $T$. Therefore, every subgroup of $\mathrm{Sym}(\Omega)$ that preserves the algebraic structure of the domain as $T^{k}/\mathrm{diag}(T^{k})$ can be written as
$T^{k}\rtimes\left(\mathrm{Out}(T)\times\mathrm{Sym}(k)\right)$
endowed with the action we just described. Therefore, a primitive permutation group $G$ is of simple diagonal type if
$T^{k}\trianglelefteq G\leq T^{k}\rtimes\left(\mathrm{Out}(T)\times\mathrm{Sym}(k)\right)\,.$
Note that, to ensure that $G$ is primitive, we need to impose that the action it induces on the $k$ direct factors of $T^{k}$ is primitive. Meanwhile $G$ is quasiprimitive, if this action is transitive.
Proof of K.
In view of Lemma 12, we can concentrate on $G=T^{k}\rtimes(\mathrm{Out}(T)\times S_{k})$, the largest primitive group of diagonal type with socle $T^{k}$ acting on $T^{k-1}$. We will build a set $A\subseteq T^{k-1}$ that cannot be separated by any of its images under the action of $G$.
First, let $B$ be any set of elements of $T$ such that
$|B|=\frac{4}{\sqrt{3}}|T|^{\frac{1}{2}},$
$1\in B$, and for every $t\in T$, $B\cap Bt$ is nonempty. (Recall that its existence is guaranteed by E.) Define
$A_{0}=B\times\ldots\times B$
as the direct product of k1k-1 copies of $B$. Note that
$|A_{0}|=|B|^{k-1}=\left(\frac{16}{3}|T|\right)^{\frac{k-1}{2}}.$
Next, let
$A_{1}=TA_{0},$
where the left multiplication by $T$ corresponds to the semiregular action of the last component of the socle. Observe that
$|A_{1}|\leq|A_{0}||T|=\left(\frac{16}{3}\right)^{\frac{k-1}{2}}|T|^{\frac{k+1}{2}}.$
Finally we enlarge $A_{1}$ one more time by
$A=A_{1}^{\mathrm{Out}(T)}.$
If $T$ is not of Lie type, then $|\mathrm{Out}(T)|\leq 4$. Meanwhile, if $T$ is of Lie type, a direct computation shows that $|\mathrm{Out}(T)|\leq\log_{2}|T|$ (see, for example, [25]). Therefore, by setting $\epsilon=1$ if $T$ is of Lie type, and $\epsilon=0$ otherwise,
$|A|$ $\leq|\mathrm{Out}(T)||A_{1}|$
$\leq 4\left(\frac{16}{3}\right)^{\frac{k-1}{2}}|T|^{\frac{k+1}{2}}\left(\frac{1}{4}\log_{2}|T|\right)^{\epsilon}$
$=4\left(\frac{16}{3}\right)^{\frac{k-1}{2}}n^{\frac{1}{2}+\frac{1}{k-1}}\left(\frac{1}{4k}\log_{2}n\right)^{\epsilon}.$
It remains to prove that the set $A$ satisfies our conditions, that is, for an arbitrary $g\in G$ we have that $A\cap A^{g}$ is nonempty. For every $g\in G$, there exist $a\in\mathrm{Sym}(k)$, $b\in\mathrm{Out}(t)$, $c\in T$ and $d\in T^{k-1}$ such that $g=abcd$. We claim that, by construction, $A$ is stabilized by $\mathrm{Sym}(k)$, by $\mathrm{Out}(T)$ and by the left semiregular action of $T$. The claim is clear for the last two items, while we shall prove it for the former. Recall that $\mathrm{Sym}(k)$ acts on $T^{k}$ by permuting components, and hence, for every point $(t_{1},\ldots,t_{k-1},1)\mathrm{diag}(T^{k})$ and for every $a\in\mathrm{Sym}(k)$,
$\mathrm{diag}(T^{k})(t_{1},t_{2},\dots,t_{k-1},1)^{a}$ $=\mathrm{diag}(T^{k})(t_{1a^{-1}},t_{2a^{-1}},\dots,t_{(k-1)x^{-1}},t_{ka^{-1}})$
$=\mathrm{diag}(T^{k})(t_{ka^{-1}}^{-1}t_{1a^{-1}},t_{ka^{-1}}^{-1}t_{2a^{-1}},\dots,t_{ka^{-1}}^{-1}t_{(k-1)a^{-1}},1).$
(For a better readability, we write the action of $\mathrm{Sym}(k)$ on the indices as a right multiplication. Moreover, to avoid a cumbersome notation, we identify $1=t_{k}$.) Observe that every direct factor of $A$ is equal, and hence the previous formula proves the claim, because
$A^{\mathrm{Sym}(k)}\subseteq TA=A.$
Hence, $A^{abc}=A$. Finally, we act with $d$ by component-wise right multiplication. By our choice of $B$, it follows that
$|A\cap A^{d}|\geq|A_{0}\cap A_{0}^{d}|\geq 1.$
Therefore, $A$ is not self-separating for $G$, which completes the proof of the upper bound for $\mathbf{m}(G)$.
The asymptotic estimate for $\mathbf{F}_{\mathcal{C}}(n)$ are obtained by restricting the attention to the groups of simple diagonal type whose socle is isomorphic to a $k$-fold Cartesian product of groups of Lie type. (Actually, groups of Lie type of characteristic $2$ are those for which the bound on $|\mathrm{Out}(T)|$ cannot be improved, see [25], and thus meet the maximum in $\mathbf{F}_{\mathcal{C}}(n)$.) ∎
We observe that it would be tempting to retrace the same proof we used for simple diagonal for every family of primitive permutation groups whose socle is regular and nonabelian. Unfortunately, this approach fails because these groups have automorphisms groups which contain an isomorphic image of the socle, and hence the resulting upper bound would exceed the degree.
Moreover, in the proof of K, our starting choice of $A_{0}$ implies that the size of $A$ depends on $k$ as an exponential function. Meanwhile, if we would have chosen $A_{0}$ to be any difference basis of the socle of appropriate size, by imposing the stability of $A$ under the action of $\mathrm{Sym}(k)$, the constant would have been factorial in $k$. Perhaps, a similar trick could extend our approach beyond the permutation groups of simple diagonal type – but there is no reason not to suspect that the trick would be specific to some limited family of socles, rather than an arbitrary semisimple group.